### v1

# Self-Ensembling Gaussian Splatting for Few-shot Novel View Synthesis

Abstract
3D Gaussian Splatting (3DGS) has demonstrated remarkable effectiveness for novel view synthesis (NVS). However, the 3DGS model tends to overfit when trained with sparse posed views, limiting its generalization capacity for broader pose variations. In this paper, we alleviate the overfitting problem by introducing a self-ensembling Gaussian Splatting (SE-GS) approach. We present two Gaussian Splatting models named the $\mathbf{\Sigma}$-model and the $\mathbf{\Delta}$-model. The $\mathbf{\Sigma}$-model serves as the primary model that generates novel-view images during inference. At the training stage, the $\mathbf{\Sigma}$-model is guided away from specific local optima by an uncertainty-aware perturbing strategy. We dynamically perturb the $\mathbf{\Delta}$-model based on the uncertainties of novel-view renderings across different training steps, resulting in diverse temporal models sampled from the Gaussian parameter space without additional training costs. The geometry of the $\mathbf{\Sigma}$-model is regularized by penalizing discrepancies between the $\mathbf{\Sigma}$-model and the temporal samples. Therefore, our SE-GS conducts an effective and efficient regularization across a large number of Gaussian Splatting models, resulting in a robust ensemble, the $\mathbf{\Sigma}$-model. Experimental results on the LLFF, Mip-NeRF360, DTU, and MVImgNet datasets show that our approach improves NVS quality with few-shot training views, outperforming existing state-of-the-art methods. The code is released at:https://github.com/sailor-z/SE-GS.
Figure 1: Qualitative results of our SE-GS and state-of-the-art approaches. The models are trained on sparse views and the images rendered from novel views are shown. As highlighted in the zoomed-in patches, our SE-GS captures finer details and produces fewer artifacts for novel views when trained on few-shot images.
1 Introduction
Novel view synthesis (NVS) is a critical task [49] in computer vision and graphics, playing a pivotal role in applications such as virtual reality [8], augmented reality [50], and 3D content generation [15, 42]. The objective of NVS is to generate photo-realistic images from previously unseen viewpoints. Typically, NVS starts by constructing a 3D representation [26, 39] from a set of existing 2D observations. In recent years, 3D Gaussian Splatting (3DGS) [20, 6, 47] has emerged as a powerful representation, integrating the advantages of both explicit [33] and implicit [26] representations. This approach enables efficient novel view generation and yields promising synthesized results with densely sampled observations that cover a wide range of viewpoints.
(a) Training
(b) Testing
Figure 2: Overfitting in 3D Gaussian Splatting with few-shot training views. 2(a) and 2(b) illustrate the performance of 3DGS on training and testing views, respectively. Each curve represents the PSNR values across different training iterations.
However, 3DGS tends to overfit the available views when only a limited number of images are provided. As shown in Fig. 2, we evaluate the 3DGS model trained on sparse images with different numbers of iterations. The performance on the training data consistently improves as the number of iterations increases, while the testing results deteriorate after 2000 iterations. Moreover, the overfitting issue becomes more noticeable with fewer training views, such as when using only 3 views. To address this problem, we introduce a new 3DGS approach in this paper, which enhances the quality of novel view synthesis with sparse training views via self-ensembling. Self-ensembling [11, 21, 27] has been witnessed in the literature as an effective strategy to mitigate overfitting, particularly in scenarios where limited data is available [45]. This technique leverages multiple predictions from the model under different conditions such as dropout [37] and data augmentation [36], forming a consensus prediction via regularization. By aggregating outputs with diverse perturbations, self-ensembling creates a more stable target for training, reducing the noise in individual predictions. Despite its success, how to utilize self-ensembling in the 3DGS pipeline remains an open problem. Therefore, we explore this gap, introducing a new 3DGS training pipeline consisting of a $\mathbf{\Sigma}$-model and a $\mathbf{\Delta}$-model.
Specifically, we present a $\mathbf{\Delta}$-model that represents a temporal sample in the Gaussian parameter space. Notably, in the scenario of sparse views, the training data lacks sufficient information to identify an optimal sample in the parameter space. As a result, the training process tends to converge to a local optimum, leading to an overfitting problem. To tackle this issue, one could create more samples by training multiple $\mathbf{\Delta}$-models, thereby increasing the likelihood of reaching a global optimum. Nevertheless, as we will demonstrate in Sec. 4, this approach incurs significant computational costs, limiting its ability to produce diverse samples. In contrast, we train a single $\mathbf{\Delta}$-model following the typical 3DGS training strategy [20], while obtaining various samples via perturbing the $\mathbf{\Delta}$-model. Rather than applying random perturbations, which often result in invalid samples, we perturb the $\mathbf{\Delta}$-model based on uncertainties derived from the training data. We store images generated by rendering the $\mathbf{\Delta}$-model from a set of pseudo views in buffers at different training iterations. These pseudo views are randomly sampled from the camera trajectories of the training views. We assess the reliability of the renderings by calculating the pixel-level uncertainties across the images within each buffer. Pixels with large uncertainty scores are identified, and the Gaussian parameters of the $\mathbf{\Delta}$-model associated with these pixels are perturbed by random noise. Therefore, our perturbation strategy generates uncertainty-aware temporal samples without introducing significant training overhead.
Building upon the $\mathbf{\Delta}$-model, we construct an ensemble of the temporal samples by proposing a $\mathbf{\Sigma}$-model. We train the $\mathbf{\Sigma}$-model over the training views without the aforementioned perturbations. To achieve self-ensembling, we regularize the geometry of the $\mathbf{\Sigma}$-model, minimizing the discrepancies between the $\mathbf{\Sigma}$-model and the perturbed $\mathbf{\Delta}$-model. The discrepancies are measured through a photometric loss between images synthesized from the pseudo views. During testing, we employ the $\mathbf{\Sigma}$-model as the primary model for novel view synthesis. Since the regularization is performed based on diverse samples in the Gaussian parameter space, the self-ensembling process enhances the robustness of the $\mathbf{\Sigma}$-model, thereby improving the generalization ability to novel views.
We conduct experiments on several datasets consisting of LLFF [25], DTU [19], Mip-NeRF360 [3], and MVImgNet [46]. The evaluation is performed with varying numbers of training views. Our SE-GS achieves the best performance across all the datasets when limited training views are used, surpassing the state-of-the-art approaches. Additionally, we provide a comprehensive analysis in which the results demonstrate the effectiveness and efficiency of our method. To the best of our knowledge, we are the first to explore the potential of the ensembling mechanism in 3DGS for few-shot novel view synthesis. We present a new self-ensembling technique that effectively mitigates the overfitting of 3DGS without requiring additional ground-truth signals or incurring significant training overhead.
2 Related Work
Radiance field modeling. Recently, Neural Radiance Fields (NeRF) [26, 41, 30, 3] has brought groundbreaking advancements to novel view synthesis. NeRF learns the density and color information of a scene or object from multi-view posed images, enabling the synthesis of realistic images from novel views via volume rendering [9]. However, most NeRF-based methods are inefficient [17] during training and inference because they require numerous MLP queries in the rendering process. This limitation makes it challenging to support tasks with real-time requirements. To this end, Kebr et al. [20] proposes an efficient alternative for radiance field modeling, which enables neural rendering via 3D Gaussian Spaltting (3DGS). 3DGS relies on explicit representations that allow for faster rendering and training [10]. Some methods [40, 5, 24, 7] explore the generalization ability of 3DGS towards novel scenes, optimizing the Gaussian parameters based on learnable cost volumes [14]. In this paper, we stick to the vanilla 3DGS rendering pipeline, focusing on the scene-level radiance field modeling.
3DGS with few-shot images. The inadequacy of training views is a primary factor leading to artifacts in novel view synthesis. Even when training views are relatively dense, some regions may still be observed from few-shot images, resulting in quality degradation in the synthesized views. To handle such an under-constrained problem, some approaches incorporate additional ground-truth signals into the 3DGS pipeline. For instance, DNGaussian [23], CoherentGS [29], and FSGS [53] utilize a monocular depth estimator [32] to predict ground-truth depth maps, enabling the 3DGS model to be trained with both photometric and depth loss. Other methods, such as [43] and [16], leverage multi-view stereo techniques [35] to generate novel-view images as ground truth. However, ground-truth data obtained from off-the-shelf methods is inevitably noisy, which potentially impacts the training of 3DGS. As noted in [48] and observed in our experiments, when the number of training views increases, the performance of these methods is sometimes worse than the vanilla 3DGS. In this context, Zhang et al. [48] proposes training multiple 3DGS models with a cross-model regularization term. This method mitigates the overfitting problem without relying on additional data. Nevertheless, training additional 3DGS models leads to significant computational costs, limiting the effectiveness of the regularization.
Ensembling. Ensembling has been witnessed as a powerful technique [52, 1, 13] in machine learning to improve model robustness and generalization by aggregating predictions from multiple models. Traditionally, ensembles [12, 31, 22] are created by training independent models and averaging their outputs or applying a voting mechanism. To improve the efficiency, some self-ensembling methods, such as temporal ensembling [21] and consistency regularization [11], leverage variations of a single model across training iterations to build an ensemble. Motivated by the success of self-ensembling, we present a Self-Ensembling Gaussian Splatting approach, in which we generate diverse samples from the Gaussian parameter space through an uncertainty-aware perturbation mechanism.
3 Method
Figure 3: Pipeline of the presented SE-GS. We tackle the scenario where only sparse posed images are available. To mitigate the overfitting of 3DGS, we incorporate a self-ensembling mechanism into the 3DGS framework. We present a $\mathbf{\Delta}$-model that represents a temporal sample in the Gaussian parameter space. This $\mathbf{\Delta}$-model is perturbed during training, which results in diverse samples. The perturbation process is carried out with the guidance of uncertainties computed from images rendered from pseudo views. The pixels in the image buffers with high uncertainties are identified and the Gaussians overlapping these pixels are perturbed, highlighted as red ellipses. The self-ensembling over the temporal samples is achieved by training a $\mathbf{\Sigma}$-model with a regularization that penalizes the discrepancies of the $\mathbf{\Sigma}$-model and the perturbed $\mathbf{\Delta}$-model. During inference, the novel view synthesis is performed using the $\mathbf{\Sigma}$-model.
3.1 Preliminaries and Motivations
3D Gaussian Splatting [20] represents a scene as a collection of Gaussians. These Gaussians are splatted onto the 2D image plane during rendering. Formally, we denote the set of $N$ Gaussians as $\{{G_{i}},i=1,2,...,N\}$. Each Gaussian $G_{i}$ is defined by $G_{i}=(\mathbf{\mu}_{i},\Sigma_{i},\mathbf{h}_{i},o_{i})$, where $\mathbf{\mu}_{i}$ is the center of the Gaussian in 3D space, $\Sigma_{i}$ is the covariance matrix, $\mathbf{h}_{i}$ represents spherical harmonics (SH) coefficients associated with the Gaussian, and $o_{i}$ indicates the opacity. Each Gaussian contributes to a 3D point $\mathbf{x}$ according to a 3D Gaussian distribution:
$G_{i}(\mathbf{x})=e^{-\frac{1}{2}(\mathbf{x}-\mathbf{\mu}_{i})^{T}\Sigma_{i}^{-1}(\mathbf{x}-\mathbf{\mu}_{i})}.$ (1)
To ensure that $\Sigma_{i}$ remains positive semi-definite throughout optimization, this matrix is decomposed into two learnable components as $\Sigma_{i}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}$, where $\mathbf{R}$ is a rotation matrix and $\mathbf{S}$ stands for a scaling matrix. For RGB rendering, the projection from 3D space to a 2D image plane is defined as:
$\Sigma_{i}^{{}^{\prime}}=\mathbf{J}\mathbf{W}\Sigma_{i}\mathbf{W}^{T}\mathbf{J}^{T},$ (2)
where $\Sigma_{i}^{{}^{\prime}}$ is the projected 2D covariance matrix, $\mathbf{W}$ represents the projection matrix, and $\mathbf{J}$ denotes the Jacobian of the affine approximation of the projection. The color of a pixel is obtained by performing alpha-blending as
$\mathbf{c}=\sum_{i=1}^{M}\mathbf{c}_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (3)
where $M$ is the number of Gaussians covering the pixel, $\mathbf{c}_{i}$ denotes the color of the Gaussian derived from the SH coefficients, and $\alpha_{i}$ is computed from the 2D covariance matrices and opacity scores. For better convergence, the Gaussian parameters are initialized based on the 3D points obtained by employing structure-from-motion (SFM) techniques [33, 34]. These parameters are optimized via gradient descent over the available posed training images.
In this paper, we focus on the scenario of sparse views, where only few-shot training images are known. We denote the 3DGS model trained on these images as $\mathcal{G}$, which represents a sample in the Gaussian parameter space. In this setting, $\mathcal{G}$ is prone to overfitting the training data and thereby getting trapped in local optima, which ultimately degrades the quality of the synthesized novel views.
To handle the overfitting problem, a promising solution [48] is to train a set of 3DGS models $\{\mathcal{G}_{1},\mathcal{G}_{2},...,\mathcal{G}_{k}\}$ that correspond to different samples. By applying regularization across this set, each model is guided away from specific local optima, facilitating a more thorough exploration of the parameter space. As we will report in Sec. 4, the performance on testing data improves as $k$ grows. However, training multiple 3DGS models incurs significant computational costs, which become unaffordable with a large $k$. This limitation therefore acts as an obstacle to scaling up the method for better performance. Drawing inspiration from the success of the regularization, we propose a novel training strategy based on self-ensembling [11, 21], which enhances 3DGS for few-shot NVS with negligible additional training overhead. The pipeline is illustrated in Fig. 3 and the details will be elaborated in this section.
3.2 Temporal Samples in 3DGS
Figure 4: Buffer update during training. For each sampled pseudo view, we dynamically update the buffer which stores the images rendered at different training steps. For instance, at training step $t_{T}$, the oldest image $\mathbf{I}_{T-S}$ in the buffer is popped, and the new image $\mathbf{I}_{T}$ is pushed into the buffer. An uncertainty map $\mathbf{U}^{t_{T}}$ is computed based on the current buffer, which is then employed for the perturbation.
In contrast to generating multiple samples by training separate 3DGS models, we dynamically create temporal samples from a single $\mathbf{\Delta}$-model during training. Specifically, we train the $\mathbf{\Delta}$-model on the available posed images, following the same optimization and density control strategies introduced in 3DGS [20]. At each training iteration, the current $\mathbf{\Delta}$-model represents a specific sample in the Gaussian parameter space based on the information contained in the training data. A temporal 3DGS model is then sampled from the neighboring region of the $\mathbf{\Delta}$-model in the parameter space by perturbing the $\mathbf{\Delta}$-model as:
$\hat{G}_{\Delta}^{t}=G_{\Delta}^{t}+\delta_{t},$ (4)
where $G_{\Delta}^{t}$ stands for a Gaussian in the $\mathbf{\Delta}$-model at the training step $t$, $\delta_{t}$ indicates the perturbations, and $\hat{G}_{\Delta}^{t}$ denotes a perturbed Gaussian in the temporal model. To achieve this, a plausible approach is to sample random noise from a normal distribution, $\delta_{t}\in\mathcal{N}(\mu_{t},\sigma_{t})$. However, since these perturbations are independent of the training data, this method tends to result in invalid samples.
Consequently, we present an uncertainty-aware perturbing strategy according to the statistics on the renderings. The approach starts by sampling pseudo views along the camera trajectories of the training views. Given two cameras sampled from the training views, the camera extrinsic of a pseudo view is computed as:
$\hat{\mathbf{R}}=\text{SLERP}(\mathbf{R}_{1},\mathbf{R}_{2},\beta),$ (5)
$\hat{\mathbf{c}}=\beta\mathbf{c}_{1}+(1-\beta)\mathbf{c}_{2},$ (6)
$\hat{\mathbf{T}}=-\hat{\mathbf{R}}\hat{\mathbf{c}},$ (7)
where $(\mathbf{R}_{1},\mathbf{R}_{2})$ represent the rotations of the sampled cameras, $(\mathbf{c}_{1},\mathbf{c}_{2})$ indicate the sampled camera centers, SLERP denotes the spherical linear interpolation [4], $\beta$ is a scalar that controls the interpolation, and $(\hat{\mathbf{R}},\hat{\mathbf{T}})$ stand for the camera parameters of the pseudo view. As illustrated in Fig. 3, for each pseudo view, we render an RGB image using the current $\mathbf{\Delta}$-model and store the renderings at different training steps in a buffer. The uncertainties of these renderings are then computed as:
$\mathbf{U}=\frac{1}{S}\sqrt{\sum_{i=1}^{S}(\mathbf{I}_{i}-\bar{\mathbf{I}})^{2}},$ (8)
where $\mathbf{U}\in\mathbb{R}^{H\times W}$ is the uncertainty map with the same resolution as the rendered image, $\mathbf{I}_{i}$ represents an image in the buffer, $\bar{\mathbf{I}}$ indicates the mean of these images, and $S$ is the buffer size. This uncertainty estimation is carried out over all sampled pseudo views in parallel. In practice, to enhance the robustness, we perform local smoothing over each uncertainty map, which is denoted as:
$\hat{\mathbf{U}}(i,j)=\frac{1}{k^{2}}\sum_{m,n}\mathbf{U}(i+m,j+n),$ (9)
where $k=5$ is the kernel size. Subsequently, we apply the uncertainty-aware perturbations to each of the Gaussians in the $\mathbf{\Delta}$-model, formulated as:
$\hat{G}_{\Delta}^{t}=G_{\Delta}^{t}+\delta_{t}h(G_{\Delta}^{t},\hat{\mathcal{U}}^{t}),$ (10)
where $\hat{\mathcal{U}}^{t}$ is a set of the computed uncertainty maps at the current training step, and $h(\cdot,\cdot)$ is an indicator function defined as:
$h(G_{\Delta}^{t},\hat{\mathcal{U}}^{t})=\begin{cases}1&\text{if }\max\limits_{(i,j)\in\mathcal{P}(G_{\Delta}^{t})}u_{ij}\geq\tau\\ 0&\text{if }\max\limits_{(i,j)\in\mathcal{P}(G_{\Delta}^{t})}u_{ij}<\tau,\end{cases}$ (11)
in which $\tau$ is a predefined threshold and $u_{ij}$ is the uncertainty score of a pixel sampled from a set of pixels $\mathcal{P}$ overlapping with the 2D splat of $G_{\Delta}^{t}$. Please refer to the supplementary material for details on defining $\tau$ and identifying $\mathcal{P}$. Therefore, the reliability of the Gaussians is connected with the uncertainties of renderings, and only the unreliable Gaussians characterized by high uncertainty scores are perturbed. In practice, we add the perturbations to the 3D points, rotations, scales, and opacities of the $\mathbf{\Delta}$-model. Particularly, since the 3D rotation is not contiguous when formulated as a 3D matrix, we perturb the 6D continuous representation [51] which is defined as:
$\hat{\mathbf{R}}^{t}_{\Delta}=f^{-1}(f(\mathbf{R}_{\Delta}^{t})+\delta^{R}_{t}h(G_{\Delta}^{t},\hat{\mathcal{U}}^{t})),\ \ \delta^{R}_{t}\in\mathbb{R}^{6}.$ (12)
Fig. 4 illustrates the dynamic updates of the buffer throughout the training process. As the buffer is updated, the uncertainty map varies accordingly, resulting in diverse temporal samples from the $\mathbf{\Delta}$-model.
Moreover, it is worth noting that the number of Gaussians in the $\mathbf{\Delta}$-model varies during training due to density control, making it challenging to directly compute the uncertainties of the Gaussian parameters across different training steps. The proposed perturbing approach can handle this issue since the statistics on 2D renderings provide a consistent and efficient assessment, regardless of the varying numbers of Gaussians.
3.3 Self-Ensembling in 3DGS
Given the temporal samples around the $\mathbf{\Delta}$-model, the next step in our pipeline is to construct an ensemble from these samples. Concretely, we train a separate 3DGS model named the $\mathbf{\Sigma}$-model on the posed images without perturbations. To prevent the $\mathbf{\Sigma}$-model from being stuck at local optima, we guide the training of the $\mathbf{\Sigma}$-model by leveraging a regularization technique [48]. The regularization is formulated as:
$\mathcal{L}_{r}=(1-\lambda){||\mathbf{I}^{t}_{\Sigma}-\mathbf{I}^{t}_{\Delta}||}_{1}+\lambda\mathcal{L}_{\text{D-SSIM}}(\mathbf{I}^{t}_{\Sigma},\mathbf{I}^{t}_{\Delta}),$ (13)
where $\lambda=0.2$ is a predefined weight, $\mathcal{L}_{\text{D-SSIM}}$ denotes a D-SSIM term, and $(\mathbf{I}^{t}_{\Sigma},\mathbf{I}^{t}_{\Delta})$ represent the images rendered from a pseudo view using the current $\mathbf{\Sigma}$-model and $\mathbf{\Delta}$-model, respectively. We also utilize a co-pruning strategy [48] to enhance the regularization. Therefore, the $\mathbf{\Sigma}$-model aggregates information from different samples in the parameter space, serving as an ensemble of 3DGS models. The final loss function during training is defined as:
$\mathcal{L}=\mathcal{L}_{\text{RGB}}+\gamma\mathcal{L}_{r},$ (14)
where $\gamma=1$ by default and $\mathcal{L}_{\text{RGB}}$ represents a photometric loss over the training views. At testing, we keep the $\mathbf{\Sigma}$-model for novel view synthesis. Notably, in contrast to CoR-GS, we derive these samples from a single $\mathbf{\Delta}$-model without additional training costs. As a result, our SE-GS can be efficiently scaled up to a large number of 3DGS models, enabling more effective regularization. Moreover, our regularization is carried out on pseudo views in a self-supervised manner, which is thus independent of additional ground truth such as monocular depth [23, 53].
[…]
3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view
Mip-NeRF [2] 16.11 22.91 24.88 0.401 0.756 0.826 0.460 0.213 0.160 0.215 0.090 0.066
DietNeRF [18] 14.94 21.75 24.28 0.370 0.717 0.801 0.496 0.248 0.183 0.240 0.105 0.073
RegNeRF [28] 19.08 23.10 24.86 0.587 0.760 0.820 0.336 0.206 0.161 0.149 0.086 0.067
FreeNeRF [44] 19.63 23.73 25.13 0.612 0.779 0.827 0.308 0.195 0.160 0.134 0.075 0.064
SparseNeRF [38] 19.86 - - 0.624 - - 0.328 - - 0.127 - -
3DGS [20] 19.22 23.80 25.44 0.649 0.814 0.860 0.229 0.125 0.096 0.120 0.066 0.051
DNGaussian [23] 19.12 - - 0.591 - - 0.294 - - 0.132 - -
FSGS [53] 20.43 24.09 25.31 0.682 0.823 0.860 0.248 0.145 0.122 0.104 0.066 0.054
CoR-GS [48] 20.45 24.49 26.06 0.712 0.837 0.874 0.196 0.115 0.089 0.101 0.060 0.046
SE-GS 20.79 24.78 26.36 0.724 0.839 0.878 0.183 0.110 0.084 0.100 0.059 0.045
Table 1: Results on LLFF with 3, 6, and 9 training views. We highlight the best, second-best, and third-best results in red, orange, and yellow, respectively.
[…]
3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view
3DGS [20] 17.67 23.69 26.80 0.804 0.894 0.941 0.158 0.086 0.050 0.112 0.058 0.031
MVSplat [7] 17.33 16.34 16.22 0.598 0.596 0.587 0.279 0.296 0.304 - - -
DNGaussian [23] 18.57 - - 0.776 - - 0.178 - - 0.110 - -
FSGS [53] 17.84 23.68 26.17 0.822 0.905 0.941 0.161 0.096 0.064 0.110 0.056 0.036
CoR-GS [48] 18.65 24.39 27.38 0.835 0.910 0.950 0.140 0.074 0.045 0.099 0.049 0.028
SE-GS 19.24 25.28 28.08 0.857 0.924 0.958 0.132 0.073 0.043 0.089 0.045 0.026
Table 2: Results on DTU with 3, 6, and 9 training views. We use red, orange, and yellow to indicate the best, second-best, and third-best results, respectively. Object masks are used for all evaluated methods to remove background when conducting the evaluation.
4 Experiments
4.1 Setup
Datasets. We conduct experiments on four datasets, i.e., LLFF [25], DTU [19], Mip-NeRF360 [3], and MVImgNet [46]. On LLFF, DTU, and Mip-NeRF360, we follow the experimental setup introduced in CoR-GS [48], using the same training/testing splits and downsampling rates. As suggested in CoR-GS, we mask the background when assessing the quality of novel view synthesis on the DTU dataset. Since LLFF, DTU, and Mip-NeRF360 offer a limited number of scenarios, we extend our experiment to a large-scale dataset, i.e., MVImgNet. We randomly sample 50 scenes from this dataset and resize the longest side of each image to 512. The Gaussian parameters for all evaluated methods are initialized based on the point clouds obtained from COLMAP [34] to ensure a fair comparison. It is worth noting that COLMAP fails on certain scenes within the DTU dataset in the setting of sparse views, where we instead randomly initialize the point clouds. Please refer to the supplementary material for more details on the setup.
Implementation details. We train our SE-GS for 10,000 iterations on the LLFF, DTU, and MVImgNet datasets, and for 30,000 iterations on Mip-NeRF360. In the uncertainty-aware perturbing mechanism, we employ 24 image buffers with a buffer size of 3 and perturb the $\mathbf{\Delta}$-model with an interval of 500 iterations. The noise $\delta_{t}$ in Eq. 10 is sampled from a standard normal distribution, i.e., $\delta_{t}\in\mathcal{N}(0,1)$. More details are provided in the supplementary material.
Method 12-view 24-view
[…]
3DGS [20] 18.52 0.523 0.415 0.167 22.80 0.708 0.276 0.096
FSGS [53] 18.80 0.531 0.418 0.163 23.28 0.715 0.274 0.093
CoR-GS [48] 19.52 0.558 0.418 0.151 23.39 0.727 0.271 0.091
SE-GS 19.91 0.596 0.400 0.143 23.74 0.745 0.265 0.086
Table 3: Results on Mip-NeRF360 with 12 and 24 training views. The first, second, and third-best results are marked in red, orange, and yellow, respectively.
[…]
5-view 7-view 10-view 5-view 7-view 10-view 5-view 7-view 10-view
3DGS [20] 26.48 29.06 31.82 0.819 0.878 0.924 0.301 0.228 0.184
FSGS [53] 26.88 29.12 31.68 0.848 0.886 0.922 0.358 0.292 0.250
CoR-GS [48] 27.62 29.77 32.09 0.842 0.889 0.928 0.302 0.235 0.190
SE-GS 27.88 30.19 32.68 0.857 0.902 0.937 0.277 0.214 0.177
Table 4: Results on MVImgNet with 5, 7, and 10 training views. Results colored in red, orange, and yellow denote the best, second-best, and third-best performances, respectively.
4.2 Quantitative results
We list the quantitative results of the evaluated approaches on the LLFF, DTU, Mip-NeRF360, and MVImgNet datasets in Table 1, Table 2, Table 3, and Table 4, respectively. The best, second-best, and third-best results in each column are highlighted in red, orange, and yellow, respectively. For MVSplat [7], we employ the pretrained model released by the authors and evaluate it on DTU. Notably, the per-scene optimized 3DGS methods significantly outperform MVSplat, highlighting the advantages of scene-level radiance field modeling for high-quality novel view synthesis. The performance of both NeRF and 3DGS methods consistently declines as the number of training views decreases, underscoring the challenge of few-shot novel view synthesis. In this context, our method surpasses all competitors across the four datasets. Note that in the scenario of sparse views, the available information in the training data is limited, making it more challenging to improve performance compared to the dense-view setting.
Specifically, on the LLFF and DTU datasets, we train the methods using 3, 6, and 9 views. Our method achieves an improvement of $0.34$ in terms of PSNR with 3 views on LLFF. The improvement is more significant on DTU, reaching a gain of $0.59$. Since the images in the Mip-NeRF360 dataset are captured from diverse viewpoints, we sample more training views, i.e., 12 and 24, following the setting used in CoR-GS. On this dataset, our SE-GS achieves the highest PSNR and SSIM, as well as the lowest LPIPS and AVGE, demonstrating superior NVS quality under the challenges of sparse views with large-scale viewpoint variations. Additionally, to further assess the robustness across diverse scenarios, we conduct an experiment on 50 scenes sampled from the MVImgNet dataset. As reported in Table 4, our method obtains consistently better results than previous state-of-the-art approaches when trained with 5, 7, and 10 views. Our SE-GS also demonstrates better compatibility with relatively dense views. For instance, the PSNR improvement increases from $0.26$ to $0.59$ as the number of training views varies from 5 to 10.
Methods like DNGaussian [23] and FSGS [53], which leverage auxiliary data terms sometimes outperform 3DGS. However, in some cases, such as 9 views on the LLFF dataset, the PSNR of FSGS becomes worse than 3DGS. Since the data acquired through off-the-shelf approaches might be unreliable, the quality of the NVS results is consequently affected. Compared with FSGS and DNGaussian, CoR-GS [48] exhibits better stability, outperforming 3DGS in more scenarios. This finding highlights the promising potential of regularization in the sparse-view setting. However, the improvement is still not consistent, as CoR-GS shows worse LPIPS than 3DGS when trained with 10 views on the MVImgNet dataset. In contrast, our SE-GS beats 3DGS in all cases, demonstrating superior stability. Our method is capable of avoiding specific local optima since the $\mathbf{\Sigma}$-model aggregates information from diverse temporal samples in the presented self-ensembling mechanism.
4.3 Qualitative results.
Figure 5: Qualitative results. The methods are trained on sparse views and the renderings of novel views are illustrated. The images are selected from the DTU and MVImgNet datasets.
Fig. 5 presents qualitative comparisons among 3DGS, CoR-GS, and our SE-GS, illustrating renderings from novel views. The models are trained on the DTU and MVImgNet datasets with 3 and 5 posed images, respectively. Our SE-GS produces fewer visual artifacts compared to the other methods, achieving better robustness against the challenge of few-shot training views. Moreover, our method captures finer details, particularly in areas with complex and repeated textures such as the pineapple and watch strap. This demonstrates the effectiveness of our self-ensembling strategy in improving both the visual quality and stability of novel view neural rendering under limited-view conditions.
(a) PSNR
(b) Training Speed
Figure 6: Comparison with CoR-GS. CoR-GS [48] is trained with different numbers of Gaussian models. The number varies from 2 to 8. PSNR on LLFF with 3 training views is utilized as the metric to evaluate the NVS quality. Iterations/s refers to the number of training iterations completed per second, indicating the training speed.
4.4 Analysis
To shed more light on the effectiveness and efficiency of our SE-GS, we develop a comprehensive analysis of the introduced self-ensembling mechanism. The analysis starts by comparing our approach with CoR-GS trained with varying numbers of Gaussian models. The detailed results on LLFF with 3 training views are shown in Fig. 6. As illustrated in Fig. 6(a), the PNSR of CoR-GS increases as more Gaussian models are trained. Since adding more Gaussian models enhances the regularization effect in CoR-GS, this observation highlights the potential of improving 3DGS in the sparse-view setting by strengthening the regularization process. The PSNR reaches the peak with 6 Gaussian models, indicating an upper bound when using a limited number of models. In contrast, our method perturbs the $\mathbf{\Delta}$-model based on the uncertainties computed from dynamically updated image buffers, which results in much more diverse samples of Gaussian models compared to CoR-GS. This leads to more effective regularization, as reflected by the better PSNR score.
Moreover, Fig. 6(b) shows the training speed measured by the number of training iterations completed per second. As witnessed in this figure, the speed of CoR-GS significantly drops as the number of Gaussian models increases. This limitation poses a challenge in scaling CoR-GS up to a large number of Gaussian models. Note that in our pipeline, only two models are trained, i.e., the $\mathbf{\Delta}$-model and $\mathbf{\Sigma}$-model. The temporal samples are generated through perturbations. The result in Fig. 6(b) shows that we achieve a comparable training speed to CoR-GS with 2 models. Only a negligible additional cost is incurred due to the perturbation process. It is worth noting that only the $\mathbf{\Sigma}$-model is retained for neural rendering during inference, so the testing speed of SE-GS is the same as that of the original 3DGS. Consequently, our method is capable of efficiently enhancing the performance of 3DGS with sparse training views.
Method Random Gradient-Aware Uncertainty-Aware
[…]
Table 5: Effectiveness of the uncertainty-aware perturbations. Random means we add random perturbations to all Gaussians in the $\mathbf{\Delta}$-model. Gradient-Aware indicates an alternative in which we perturb the Gaussians based on gradients. All methods are trained on LLFF with 3 views for 2000 iterations and the PSNR is utilized as the metric.
Furthermore, we assess the effectiveness of our uncertainty-aware perturbing strategy, comparing it with two alternatives. Specifically, as we introduced in Sec. 3, the straightforward method perturbing the $\mathbf{\Delta}$-model is to add random noise to all Gaussians in the model. We denote this approach as Random. Building upon this baseline, an alternative approach, referred to as Gradient-Aware, identifies unreliable Gaussians by analyzing their gradients. Specifically, we replace the indicator function in Eq. 11 with one based on gradient magnitude. Gaussians with gradients exceeding a threshold are perturbed. In practice, we adopt the same threshold as used in density control. We train these two alternatives and our approaches on LLFF with 3 views for 2000 iterations, which challenges the methods in terms of convergence speed. As listed in Table 5, Random performs the worst since it leads to many invalid Gaussian samples, negatively impacting convergence speed. Note that our uncertainty-aware strategy aggregates information from multiple training steps, while the gradient-aware method can only utilize the gradients at the current training step. The improvement in PSNR shown in Table 5 demonstrates the superior effectiveness of our approach.
5 Conclusion
In this paper, we have introduced a new self-ensembling mechanism that enhances the novel view synthesis of 3DGS with few-shot training images. We tackle the challenge of overfitting by training an ensemble model named $\mathbf{\Sigma}$-model with the guidance of diverse temporal samples from the Gaussian parameter space. We introduce an effective sampling strategy that perturbs a $\mathbf{\Delta}$-model based on the uncertainties computed from dynamically updated buffers of pseudo-view renderings. We have conducted experiments on LLFF, DTU, Mip-NeRF360, and MVImgNet, as well as a comprehensive analysis of our approach. The experimental results have demonstrated the effectiveness, stability, and efficiency of our SE-GS.
Acknowledgment. This work was funded in part by the Swiss National Science Foundation via the Sinergia grant CRSII5-180359 and the Swiss Innovation Agency (Innosuisse) via the BRIDGE Discovery grant 40B2-0_194729.

### v2

@@
 3D Gaussian Splatting (3DGS) has demonstrated remarkable effectiveness for novel view synthesis (NVS).
-However, the 3DGS model tends to overfit when trained with sparse posed views, limiting its generalization capacity for broader pose variations.
-In this paper, we alleviate the overfitting problem by introducing a self-ensembling Gaussian Splatting (SE-GS) approach.
-We present two Gaussian Splatting models named the $\mathbf{\Sigma}$-model and the $\mathbf{\Delta}$-model.
-The $\mathbf{\Sigma}$-model serves as the primary model that generates novel-view images during inference.
-At the training stage, the $\mathbf{\Sigma}$-model is guided away from specific local optima by an uncertainty-aware perturbing strategy.
-We dynamically perturb the $\mathbf{\Delta}$-model based on the uncertainties of novel-view renderings across different training steps, resulting in diverse temporal models sampled from the Gaussian parameter space without additional training costs.
-The geometry of the $\mathbf{\Sigma}$-model is regularized by penalizing discrepancies between the $\mathbf{\Sigma}$-model and the temporal samples.
-Therefore, our SE-GS conducts an effective and efficient regularization across a large number of Gaussian Splatting models, resulting in a robust ensemble, the $\mathbf{\Sigma}$-model.
-Experimental results on the LLFF, Mip-NeRF360, DTU, and MVImgNet datasets show that our approach improves NVS quality with few-shot training views, outperforming existing state-of-the-art methods.
-The code is released at:https://github.com/sailor-z/SE-GS.
+However, the 3DGS model tends to overfit when trained with sparse posed views, limiting its generalization ability to novel views.
+In this paper, we alleviate the overfitting problem, presenting a Self-Ensembling Gaussian Splatting (SE-GS) approach.
+Our method encompasses a $\mathbf{\Sigma}$-model and a $\mathbf{\Delta}$-model.
+The $\mathbf{\Sigma}$-model serves as an ensemble of 3DGS models that generates novel-view images during inference.
+We achieve the self-ensembling by introducing an uncertainty-aware perturbation strategy at the training state.
+We complement the $\mathbf{\Sigma}$-model with the $\mathbf{\Delta}$-model, which is dynamically perturbed based on the uncertainties of novel-view renderings across different training steps.
+The perturbation yields diverse temporal samples in the Gaussian parameter space without additional training costs.
+The geometry of the $\mathbf{\Sigma}$-model is regularized by penalizing discrepancies between the $\mathbf{\Sigma}$-model and these temporal samples.
+Therefore, our SE-GS conducts an effective and efficient regularization across a large number of 3DGS models, resulting in a robust ensemble, the $\mathbf{\Sigma}$-model.
+Our experimental results on the LLFF, Mip-NeRF360, DTU, and MVImgNet datasets show that our approach improves NVS quality with few-shot training views, outperforming existing state-of-the-art methods.
+The code is released at:project page.
 Figure 1: Qualitative results of our SE-GS and state-of-the-art approaches.
@@
 This approach enables efficient novel view generation and yields promising synthesized results with densely sampled observations that cover a wide range of viewpoints.
+However, 3DGS tends to overfit the available views when only a limited number of images are provided.
+To illustrate this issue, in Fig.
+2, we evaluate the 3DGS model trained on sparse images with different numbers of iterations.
+The performance on the training data consistently improves as the number of iterations increases, while the testing results deteriorate after 2000 iterations.
+Moreover, the overfitting problem becomes more noticeable with fewer training views, such as when using only 3 views.
 (a) Training
 (b) Testing
-Figure 2: Overfitting in 3D Gaussian Splatting with few-shot training views.
+Figure 2: Overfitting in 3D Gaussian Splatting with sparse training views.
 2(a) and 2(b) illustrate the performance of 3DGS on training and testing views, respectively.
 Each curve represents the PSNR values across different training iterations.
-However, 3DGS tends to overfit the available views when only a limited number of images are provided.
-As shown in Fig.
-2, we evaluate the 3DGS model trained on sparse images with different numbers of iterations.
-The performance on the training data consistently improves as the number of iterations increases, while the testing results deteriorate after 2000 iterations.
-Moreover, the overfitting issue becomes more noticeable with fewer training views, such as when using only 3 views.
-To address this problem, we introduce a new 3DGS approach in this paper, which enhances the quality of novel view synthesis with sparse training views via self-ensembling.
-Self-ensembling [11, 21, 27] has been witnessed in the literature as an effective strategy to mitigate overfitting, particularly in scenarios where limited data is available [45].
-This technique leverages multiple predictions from the model under different conditions such as dropout [37] and data augmentation [36], forming a consensus prediction via regularization.
-By aggregating outputs with diverse perturbations, self-ensembling creates a more stable target for training, reducing the noise in individual predictions.
-Despite its success, how to utilize self-ensembling in the 3DGS pipeline remains an open problem.
-Therefore, we explore this gap, introducing a new 3DGS training pipeline consisting of a $\mathbf{\Sigma}$-model and a $\mathbf{\Delta}$-model.
-Specifically, we present a $\mathbf{\Delta}$-model that represents a temporal sample in the Gaussian parameter space.
-Notably, in the scenario of sparse views, the training data lacks sufficient information to identify an optimal sample in the parameter space.
-As a result, the training process tends to converge to a local optimum, leading to an overfitting problem.
-To tackle this issue, one could create more samples by training multiple $\mathbf{\Delta}$-models, thereby increasing the likelihood of reaching a global optimum.
+To mitigate overfitting, self-ensembling [11, 22, 29] has been highlighted as an effective strategy in tasks such as detection [48] and segmentation [25].
+Nevertheless, its use for NVS remains unstudied, and how to exploit it in a 3DGS formalism is an open question.
+Therefore, in this paper, we bridge this gap, introducing a new 3DGS method that enhances the quality of novel view synthesis with sparse training views via self-ensembling.
+Our method consists of two models, a $\mathbf{\Sigma}$-model and a $\mathbf{\Delta}$-model.
+The $\mathbf{\Sigma}$-model constitutes an ensemble of 3DGS models, which is used to generate novel views during inference.
+A straightforward way to achieve ensembling would be to utilize multiple 3DGS models that correspond to different samples in the Gaussian parameter space, thereby increasing the likelihood of reaching a more robust solution.
 Nevertheless, as we will demonstrate in Sec.
-4, this approach incurs significant computational costs, limiting its ability to produce diverse samples.
-In contrast, we train a single $\mathbf{\Delta}$-model following the typical 3DGS training strategy [20], while obtaining various samples via perturbing the $\mathbf{\Delta}$-model.
-Rather than applying random perturbations, which often result in invalid samples, we perturb the $\mathbf{\Delta}$-model based on uncertainties derived from the training data.
-We store images generated by rendering the $\mathbf{\Delta}$-model from a set of pseudo views in buffers at different training iterations.
-These pseudo views are randomly sampled from the camera trajectories of the training views.
-We assess the reliability of the renderings by calculating the pixel-level uncertainties across the images within each buffer.
-Pixels with large uncertainty scores are identified, and the Gaussian parameters of the $\mathbf{\Delta}$-model associated with these pixels are perturbed by random noise.
-Therefore, our perturbation strategy generates uncertainty-aware temporal samples without introducing significant training overhead.
-Building upon the $\mathbf{\Delta}$-model, we construct an ensemble of the temporal samples by proposing a $\mathbf{\Sigma}$-model.
-We train the $\mathbf{\Sigma}$-model over the training views without the aforementioned perturbations.
-To achieve self-ensembling, we regularize the geometry of the $\mathbf{\Sigma}$-model, minimizing the discrepancies between the $\mathbf{\Sigma}$-model and the perturbed $\mathbf{\Delta}$-model.
-The discrepancies are measured through a photometric loss between images synthesized from the pseudo views.
-During testing, we employ the $\mathbf{\Sigma}$-model as the primary model for novel view synthesis.
-Since the regularization is performed based on diverse samples in the Gaussian parameter space, the self-ensembling process enhances the robustness of the $\mathbf{\Sigma}$-model, thereby improving the generalization ability to novel views.
-We conduct experiments on several datasets consisting of LLFF [25], DTU [19], Mip-NeRF360 [3], and MVImgNet [46].
-The evaluation is performed with varying numbers of training views.
-Our SE-GS achieves the best performance across all the datasets when limited training views are used, surpassing the state-of-the-art approaches.
-Additionally, we provide a comprehensive analysis in which the results demonstrate the effectiveness and efficiency of our method.
-To the best of our knowledge, we are the first to explore the potential of the ensembling mechanism in 3DGS for few-shot novel view synthesis.
-We present a new self-ensembling technique that effectively mitigates the overfitting of 3DGS without requiring additional ground-truth signals or incurring significant training overhead.
+4, the computational cost of training multiple 3DGS models limits the ability to produce sufficiently diverse samples.
+By contrast, we complement the $\mathbf{\Sigma}$-model with a single $\mathbf{\Delta}$-model and obtain diverse samples by perturbing the $\mathbf{\Delta}$-model.
+Specifically, instead of applying random perturbation, which often results in invalid samples, we perturb the $\mathbf{\Delta}$-model based on uncertainties derived from the training data.
+We store images generated by rendering the $\mathbf{\Delta}$-model from a set of pseudo views at different training iterations in buffers and calculate pixel-level uncertainties across images within each buffer.
+We then add random noise to the Gaussian parameters of the $\mathbf{\Delta}$-model associated with pixels that have high uncertainty scores.
+Consequently, our perturbation strategy generates uncertainty-aware samples without introducing significant training overhead.
+Building upon the uncertainty-aware samples derived from the $\mathbf{\Delta}$-model, we construct an ensemble by leveraging a self-ensembling mechanism.
+We train the $\mathbf{\Sigma}$-model over the training views without the aforementioned perturbation and regularize the geometry of the $\mathbf{\Sigma}$-model by minimizing its discrepancies with the perturbed $\mathbf{\Delta}$-model.
+These discrepancies are measured via a photometric loss between images synthesized from the pseudo views.
+Since the regularization is performed based on diverse samples in the Gaussian parameter space, the self-ensembling process enhances the robustness of the $\mathbf{\Sigma}$-model, thereby improving its generalization ability to novel views.
+Moreover, our self-ensembling technique is independent of any additional ground-truth signals as the regularization is carried out in a self-supervised manner.
+We conduct experiments on multiple datasets, including LLFF [27], DTU [20], Mip-NeRF360 [3], and MVImgNet [45], with limited numbers of training views.
+Our SE-GS achieves the best performance across all datasets in the sparse-view setting, surpassing the state-of-the-art approaches.
+Furthermore, we perform a comprehensive analysis of our method to demonstrate its effectiveness and efficiency.
+To the best of our knowledge, we are the first to explore the potential of the self-ensembling mechanism in 3DGS for few-shot novel view synthesis.
+The code will be publicly available upon acceptance.
 2 Related Work
 Radiance field modeling.
-Recently, Neural Radiance Fields (NeRF) [26, 41, 30, 3] has brought groundbreaking advancements to novel view synthesis.
+Recently, Neural Radiance Field (NeRF) [28, 41, 32, 3] has brought groundbreaking advancements to novel view synthesis.
 NeRF learns the density and color information of a scene or object from multi-view posed images, enabling the synthesis of realistic images from novel views via volume rendering [9].
@@
 This limitation makes it challenging to support tasks with real-time requirements.
-To this end, Kebr et al.
-[20] proposes an efficient alternative for radiance field modeling, which enables neural rendering via 3D Gaussian Spaltting (3DGS).
-3DGS relies on explicit representations that allow for faster rendering and training [10].
-Some methods [40, 5, 24, 7] explore the generalization ability of 3DGS towards novel scenes, optimizing the Gaussian parameters based on learnable cost volumes [14].
+In this context, 3D Gaussian Splatting [21] is developed as an efficient alternative for radiance field modeling.
+3DGS relies on explicit representations [35] that allow for faster rendering and training [10].
+While initially introduced as a scene-specific model, some 3DGS variants [40, 5, 26, 7] explore the generalization ability of 3DGS towards novel scenes, optimizing the Gaussian parameters based on learnable cost volumes [15].
 In this paper, we stick to the vanilla 3DGS rendering pipeline, focusing on the scene-level radiance field modeling.
 3DGS with few-shot images.
-The inadequacy of training views is a primary factor leading to artifacts in novel view synthesis.
+The insufficiency of training views is a primary factor leading to artifacts in novel view synthesis.
 Even when training views are relatively dense, some regions may still be observed from few-shot images, resulting in quality degradation in the synthesized views.
 To handle such an under-constrained problem, some approaches incorporate additional ground-truth signals into the 3DGS pipeline.
-For instance, DNGaussian [23], CoherentGS [29], and FSGS [53] utilize a monocular depth estimator [32] to predict ground-truth depth maps, enabling the 3DGS model to be trained with both photometric and depth loss.
+For instance, DNGaussian [24], CoherentGS [31], and FSGS [53] utilize a monocular depth estimator [34] to predict depth maps, enabling the 3DGS model to be trained with both a photometric and a depth loss.
 Other methods, such as [43] and [17], leverage multi-view stereo techniques [37] to generate novel-view images as ground truth.
@@
 As noted in [47] and observed in our experiments, when the number of training views increases, the performance of these methods is sometimes worse than the vanilla 3DGS.
-In this context, Zhang et al.
-[48] proposes training multiple 3DGS models with a cross-model regularization term.
-This method mitigates the overfitting problem without relying on additional data.
-Nevertheless, training additional 3DGS models leads to significant computational costs, limiting the effectiveness of the regularization.
+To overcome this problem, Zhang et al.
+[47] propose training multiple 3DGS models with a cross-model regularization term.
+However, training additional 3DGS models incurs a significant computational cost, making it impractical to scale up to a large number of 3DGS models for stronger regularization.
 Ensembling.
-Ensembling has been witnessed as a powerful technique [52, 1, 13] in machine learning to improve model robustness and generalization by aggregating predictions from multiple models.
+Ensembling has been evidenced as a powerful technique in machine learning [52, 1, 14] to improve model robustness and generalization by aggregating predictions from multiple models.
 Traditionally, ensembles [13, 33, 23] are created by training independent models and averaging their outputs or applying a voting mechanism.
 To improve the efficiency, some self-ensembling methods, such as temporal ensembling [22] and consistency regularization [11], leverage variations of a single model across training iterations to build an ensemble.
-Motivated by the success of self-ensembling, we present a Self-Ensembling Gaussian Splatting approach, in which we generate diverse samples from the Gaussian parameter space through an uncertainty-aware perturbation mechanism.
+Motivated by the success of self-ensembling, we present the first Self-Ensembling Gaussian Splatting approach, in which we generate diverse samples in the Gaussian parameter space through an uncertainty-aware perturbation mechanism.
 3 Method
 Figure 3: Pipeline of the presented SE-GS.
-We tackle the scenario where only sparse posed images are available.
-To mitigate the overfitting of 3DGS, we incorporate a self-ensembling mechanism into the 3DGS framework.
-We present a $\mathbf{\Delta}$-model that represents a temporal sample in the Gaussian parameter space.
-This $\mathbf{\Delta}$-model is perturbed during training, which results in diverse samples.
-The perturbation process is carried out with the guidance of uncertainties computed from images rendered from pseudo views.
-The pixels in the image buffers with high uncertainties are identified and the Gaussians overlapping these pixels are perturbed, highlighted as red ellipses.
-The self-ensembling over the temporal samples is achieved by training a $\mathbf{\Sigma}$-model with a regularization that penalizes the discrepancies of the $\mathbf{\Sigma}$-model and the perturbed $\mathbf{\Delta}$-model.
-During inference, the novel view synthesis is performed using the $\mathbf{\Sigma}$-model.
+We tackle the overfitting problem in sparse-view scenarios by incorporating a self-ensembling mechanism into the 3DGS framework.
+Our self-ensembling builds upon diverse samples in the Gaussian parameter space.
+These samples are derived from the $\mathbf{\Delta}$-model via an uncertainty-aware perturbation strategy.
+Specifically, we store images rendered from pseudo views at different training steps in buffers, from which we compute pixel-level uncertainties.
+We then perturb the Gaussians overlapping the pixels with high uncertainties, as highlighted as red ellipses.
+Self-ensembling over the resulting samples is achieved by training a $\mathbf{\Sigma}$-model with a regularization that penalizes the discrepancies of the $\mathbf{\Sigma}$-model and the perturbed $\mathbf{\Delta}$-model.
+During inference, novel view synthesis is performed using the $\mathbf{\Sigma}$-model.
 3.1 Preliminaries and Motivations
@@
 Formally, we denote the set of $N$ Gaussians as $\{{G_{i}},i=1,2,...,N\}$.
-Each Gaussian $G_{i}$ is defined by $G_{i}=(\mathbf{\mu}_{i},\Sigma_{i},\mathbf{h}_{i},o_{i})$, where $\mathbf{\mu}_{i}$ is the center of the Gaussian in 3D space, $\Sigma_{i}$ is the covariance matrix, $\mathbf{h}_{i}$ represents spherical harmonics (SH) coefficients associated with the Gaussian, and $o_{i}$ indicates the opacity.
-Each Gaussian contributes to a 3D point $\mathbf{x}$ according to a 3D Gaussian distribution:
-$G_{i}(\mathbf{x})=e^{-\frac{1}{2}(\mathbf{x}-\mathbf{\mu}_{i})^{T}\Sigma_{i}^{-1}(\mathbf{x}-\mathbf{\mu}_{i})}.$ (1)
-To ensure that $\Sigma_{i}$ remains positive semi-definite throughout optimization, this matrix is decomposed into two learnable components as $\Sigma_{i}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}$, where $\mathbf{R}$ is a rotation matrix and $\mathbf{S}$ stands for a scaling matrix.
-For RGB rendering, the projection from 3D space to a 2D image plane is defined as:
+Each Gaussian $G_{i}$ is defined as $G_{i}=(\mathbf{\mu}_{i},\Sigma_{i},\mathbf{h}_{i},o_{i})$, where $\mathbf{\mu}_{i}$ is the 3D position, $\Sigma_{i}$ is the covariance matrix, $\mathbf{h}_{i}$ represents spherical harmonics (SH) coefficients associated with the Gaussian, and $o_{i}$ indicates opacity.
+Each Gaussian contributes to a 3D point $\mathbf{x}$ according to the 3D Gaussian distribution
+$\mathcal{N}_{3d}^{i}(\mathbf{x})=e^{-\frac{1}{2}(\mathbf{x}-\mathbf{\mu}_{i})^{T}\Sigma_{i}^{-1}(\mathbf{x}-\mathbf{\mu}_{i})}.$ (1)
+The parameters of the Gaussians are then learned from the input views.
+To ensure that $\Sigma_{i}$ remains positive semi-definite throughout optimization, it is decomposed into two learnable components as $\Sigma_{i}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}$, where $\mathbf{R}$ is a rotation matrix and $\mathbf{S}$ stands for a scaling matrix.
+For RGB rendering, the projection from 3D space to a 2D image plane relies on the projection matrix $\mathbf{W}$ to compute the projected 2D covariance matrix
 $\Sigma_{i}^{{}^{\prime}}=\mathbf{J}\mathbf{W}\Sigma_{i}\mathbf{W}^{T}\mathbf{J}^{T},$ (2)
-where $\Sigma_{i}^{{}^{\prime}}$ is the projected 2D covariance matrix, $\mathbf{W}$ represents the projection matrix, and $\mathbf{J}$ denotes the Jacobian of the affine approximation of the projection.
-The color of a pixel is obtained by performing alpha-blending as
+where $\mathbf{J}$ denotes the Jacobian of the affine approximation of the projection. The color of a pixel is obtained by performing alpha-blending as
 $\mathbf{c}=\sum_{i=1}^{M}\mathbf{c}_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (3)
 where $M$ is the number of Gaussians covering the pixel, $\mathbf{c}_{i}$ denotes the color of the Gaussian derived from the SH coefficients, and $\alpha_{i}$ is computed from the 2D covariance matrices and opacity scores.
-For better convergence, the Gaussian parameters are initialized based on the 3D points obtained by employing structure-from-motion (SFM) techniques [33, 34].
-These parameters are optimized via gradient descent over the available posed training images.
-In this paper, we focus on the scenario of sparse views, where only few-shot training images are known.
+For better convergence, the Gaussian parameters are initialized based on 3D points obtained using structure-from-motion (SFM) techniques [35, 36].
+These parameters are optimized via gradient descent with a photometric loss [21] where the posed training images serve as ground truth.
+In this paper, we focus on sparse-view scenarios, where only few-shot training images are provided.
 We denote the 3DGS model trained on these images as $\mathcal{G}$, which represents a sample in the Gaussian parameter space.
-In this setting, $\mathcal{G}$ is prone to overfitting the training data and thereby getting trapped in local optima, which ultimately degrades the quality of the synthesized novel views.
-To handle the overfitting problem, a promising solution [48] is to train a set of 3DGS models $\{\mathcal{G}_{1},\mathcal{G}_{2},...,\mathcal{G}_{k}\}$ that correspond to different samples.
+In this setting, $\mathcal{G}$ is prone to overfitting the training data and thereby getting trapped in a suboptimal solution, which ultimately degrades the quality of the synthesized novel views.
+To handle the overfitting problem, a promising solution is to train a set of 3DGS models $\{\mathcal{G}_{1},\mathcal{G}_{2},...,\mathcal{G}_{k}\}$ that correspond to different samples.
 By applying regularization across this set, each model is guided away from specific local optima, facilitating a more thorough exploration of the parameter space.
@@
 4, the performance on testing data improves as $k$ grows.
-However, training multiple 3DGS models incurs significant computational costs, which become unaffordable with a large $k$.
-This limitation therefore acts as an obstacle to scaling up the method for better performance.
-Drawing inspiration from the success of the regularization, we propose a novel training strategy based on self-ensembling [11, 21], which enhances 3DGS for few-shot NVS with negligible additional training overhead.
+However, training multiple 3DGS models incurs a significant computational cost, which becomes unaffordable with a large $k$.
+This limitation prevents the method from scaling up for better performance.
+Consequently, we propose a novel training strategy based on self-ensembling [11, 22], which enhances 3DGS for few-shot NVS with negligible additional training overhead.
 The pipeline is illustrated in Fig.
-3 and the details will be elaborated in this section.
-3.2 Temporal Samples in 3DGS
+3 and the details will be elaborated in the remainder of this section.
+3.2 Uncertainty-Aware Perturbation
 Figure 4: Buffer update during training.
-For each sampled pseudo view, we dynamically update the buffer which stores the images rendered at different training steps.
+For each sampled pseudo view, we dynamically update the buffer that stores the images rendered at different training steps.
 For instance, at training step $t_{T}$, the oldest image $\mathbf{I}_{T-S}$ in the buffer is popped, and the new image $\mathbf{I}_{T}$ is pushed into the buffer.
-An uncertainty map $\mathbf{U}^{t_{T}}$ is computed based on the current buffer, which is then employed for the perturbation.
+An uncertainty map $\mathbf{U}^{t_{T}}$ is computed based on the current buffer, which is then employed to compute a perturbation that results in a new 3DGS model.
 In contrast to generating multiple samples by training separate 3DGS models, we dynamically create temporal samples from a single $\mathbf{\Delta}$-model during training.
+Note that, in this paper, temporal refers to the dynamic nature of samples that vary across training iterations.
 Specifically, we train the $\mathbf{\Delta}$-model on the available posed images, following the same optimization and density control strategies introduced in 3DGS [21].
 At each training iteration, the current $\mathbf{\Delta}$-model represents a specific sample in the Gaussian parameter space based on the information contained in the training data.
-A temporal 3DGS model is then sampled from the neighboring region of the $\mathbf{\Delta}$-model in the parameter space by perturbing the $\mathbf{\Delta}$-model as:
+A temporal 3DGS model is then sampled from the neighboring region of the current $\mathbf{\Delta}$-model in the parameter space by perturbing the $\mathbf{\Delta}$-model as
 $\hat{G}_{\Delta}^{t}=G_{\Delta}^{t}+\delta_{t},$ (4)
-where $G_{\Delta}^{t}$ stands for a Gaussian in the $\mathbf{\Delta}$-model at the training step $t$, $\delta_{t}$ indicates the perturbations, and $\hat{G}_{\Delta}^{t}$ denotes a perturbed Gaussian in the temporal model.
-To achieve this, a plausible approach is to sample random noise from a normal distribution, $\delta_{t}\in\mathcal{N}(\mu_{t},\sigma_{t})$.
-However, since these perturbations are independent of the training data, this method tends to result in invalid samples.
-Consequently, we present an uncertainty-aware perturbing strategy according to the statistics on the renderings.
-The approach starts by sampling pseudo views along the camera trajectories of the training views.
-Given two cameras sampled from the training views, the camera extrinsic of a pseudo view is computed as:
+where $G_{\Delta}^{t}$ stands for a Gaussian in the $\mathbf{\Delta}$-model at training step $t$, $\delta_{t}$ indicates the perturbation, and $\hat{G}_{\Delta}^{t}$ denotes a perturbed Gaussian in the temporal model.
+To achieve the perturbation, a naive approach is to sample random noise from a normal distribution, i.e., $\delta_{t}\in\mathcal{N}(\mu_{t},\sigma_{t}^{2})$.
+However, such perturbation is independent of the training data, and this method thus tends to yield invalid samples.
+Therefore, we present an uncertainty-aware perturbation strategy, leveraging the statistics of renderings.
+Our approach starts by creating $M$ pseudo views through interpolation between the training views.
+Specifically, given two cameras sampled from the training views, the camera extrinsics of a pseudo view are computed as
 $\hat{\mathbf{R}}=\text{SLERP}(\mathbf{R}_{1},\mathbf{R}_{2},\beta),$ (5)
@@
 3, for each pseudo view, we render an RGB image using the current $\mathbf{\Delta}$-model and store the renderings at different training steps in a buffer.
-The uncertainties of these renderings are then computed as:
-$\mathbf{U}=\frac{1}{S}\sqrt{\sum_{i=1}^{S}(\mathbf{I}_{i}-\bar{\mathbf{I}})^{2}},$ (8)
-where $\mathbf{U}\in\mathbb{R}^{H\times W}$ is the uncertainty map with the same resolution as the rendered image, $\mathbf{I}_{i}$ represents an image in the buffer, $\bar{\mathbf{I}}$ indicates the mean of these images, and $S$ is the buffer size.
-This uncertainty estimation is carried out over all sampled pseudo views in parallel.
-In practice, to enhance the robustness, we perform local smoothing over each uncertainty map, which is denoted as:
+We then compute a pixel-wise uncertainty map
+$\mathbf{U}=\sqrt{\frac{1}{S}\sum_{i=1}^{S}(\mathbf{I}_{i}-\bar{\mathbf{I}})^{2}},$ (8)
+of the same resolution as the rendered image, i.e., $\mathbf{U}\in\mathbb{R}^{H\times W}$, where $\mathbf{I}_{i}$ represents an image in the buffer, $\bar{\mathbf{I}}$ indicates the mean of these images, and $S$ is the buffer size.
+This uncertainty estimation is carried out over all $M$ sampled pseudo views in parallel.
+In practice, to enhance robustness, we perform local smoothing over each uncertainty map, which yields
 $\hat{\mathbf{U}}(i,j)=\frac{1}{k^{2}}\sum_{m,n}\mathbf{U}(i+m,j+n),$ (9)
-where $k=5$ is the kernel size. Subsequently, we apply the uncertainty-aware perturbations to each of the Gaussians in the $\mathbf{\Delta}$-model, formulated as:
+where $k=5$ is the kernel size. Subsequently, we apply an uncertainty-aware perturbation to each of the Gaussians in the $\mathbf{\Delta}$-model as
 $\hat{G}_{\Delta}^{t}=G_{\Delta}^{t}+\delta_{t}h(G_{\Delta}^{t},\hat{\mathcal{U}}^{t}),$ (10)
-where $\hat{\mathcal{U}}^{t}$ is a set of the computed uncertainty maps at the current training step, and $h(\cdot,\cdot)$ is an indicator function defined as:
+where $\hat{\mathcal{U}}^{t}$ is the set of the uncertainty maps computed from the buffers at the current training step, and $h(\cdot,\cdot)$ is an indicator function defined as
 $h(G_{\Delta}^{t},\hat{\mathcal{U}}^{t})=\begin{cases}1&\text{if }\max\limits_{(i,j)\in\mathcal{P}(G_{\Delta}^{t})}u_{ij}\geq\tau\\ 0&\text{if }\max\limits_{(i,j)\in\mathcal{P}(G_{\Delta}^{t})}u_{ij}<\tau,\end{cases}$ (11)
-in which $\tau$ is a predefined threshold and $u_{ij}$ is the uncertainty score of a pixel sampled from a set of pixels $\mathcal{P}$ overlapping with the 2D splat of $G_{\Delta}^{t}$.
+with $\tau$ a predefined threshold, and $u_{ij}$ the uncertainty score of a pixel sampled from the set of pixels $\mathcal{P}$ overlapping with the 2D splats of $G_{\Delta}^{t}$.
 Please refer to the supplementary material for details on defining $\tau$ and identifying $\mathcal{P}$.
-Therefore, the reliability of the Gaussians is connected with the uncertainties of renderings, and only the unreliable Gaussians characterized by high uncertainty scores are perturbed.
-In practice, we add the perturbations to the 3D points, rotations, scales, and opacities of the $\mathbf{\Delta}$-model.
-Particularly, since the 3D rotation is not contiguous when formulated as a 3D matrix, we perturb the 6D continuous representation [51] which is defined as:
-$\hat{\mathbf{R}}^{t}_{\Delta}=f^{-1}(f(\mathbf{R}_{\Delta}^{t})+\delta^{R}_{t}h(G_{\Delta}^{t},\hat{\mathcal{U}}^{t})),\ \ \delta^{R}_{t}\in\mathbb{R}^{6}.$ (12)
+In short, the reliability of each Gaussian is connected with the uncertainties of the renderings, and only the unreliable Gaussians, characterized by high uncertainty scores, are perturbed.
+In practice, we add the perturbation to the 3D positions, 3D rotations, scales, and opacities of the $\mathbf{\Delta}$-model.
+In particular, since 3D rotation is not continuous when expressed as a 3D matrix, we perturb the 6D continuous representation [51] of the rotation, which yields
+$\hat{\mathbf{R}}^{t}_{\Delta}=f^{-1}(f(\mathbf{R}_{\Delta}^{t})+\delta^{R}_{t}h(G_{\Delta}^{t},\hat{\mathcal{U}}^{t})),\ \ \delta^{R}_{t}\in\mathbb{R}^{6},$ (12)
+where $f(\cdot)$ denotes the mapping from rotation matrix to 6D representation.
 Fig.
@@
 As the buffer is updated, the uncertainty map varies accordingly, resulting in diverse temporal samples from the $\mathbf{\Delta}$-model.
-Moreover, it is worth noting that the number of Gaussians in the $\mathbf{\Delta}$-model varies during training due to density control, making it challenging to directly compute the uncertainties of the Gaussian parameters across different training steps.
-The proposed perturbing approach can handle this issue since the statistics on 2D renderings provide a consistent and efficient assessment, regardless of the varying numbers of Gaussians.
+It is worth noting that the number of Gaussians in the $\mathbf{\Delta}$-model varies during training due to density control, making it challenging to compute uncertainties directly in the Gaussian parameter space across different training steps.
+Our approach naturally handles this challenge as the statistics on 2D renderings provide a consistent ground to measure uncertainties, regardless of the varying number of Gaussians.
 3.3 Self-Ensembling in 3DGS
 Given the temporal samples around the $\mathbf{\Delta}$-model, the next step in our pipeline is to construct an ensemble from these samples.
-Concretely, we train a separate 3DGS model named the $\mathbf{\Sigma}$-model on the posed images without perturbations.
-To prevent the $\mathbf{\Sigma}$-model from being stuck at local optima, we guide the training of the $\mathbf{\Sigma}$-model by leveraging a regularization technique [48].
-The regularization is formulated as:
+Concretely, we train a separate 3DGS model, named the $\mathbf{\Sigma}$-model, on the posed images without perturbations.
+To prevent the $\mathbf{\Sigma}$-model from being trapped at a suboptimal solution, we guide the training of the $\mathbf{\Sigma}$-model with a regularization [47] formulated as
 $\mathcal{L}_{r}=(1-\lambda){||\mathbf{I}^{t}_{\Sigma}-\mathbf{I}^{t}_{\Delta}||}_{1}+\lambda\mathcal{L}_{\text{D-SSIM}}(\mathbf{I}^{t}_{\Sigma},\mathbf{I}^{t}_{\Delta}),$ (13)
-where $\lambda=0.2$ is a predefined weight, $\mathcal{L}_{\text{D-SSIM}}$ denotes a D-SSIM term, and $(\mathbf{I}^{t}_{\Sigma},\mathbf{I}^{t}_{\Delta})$ represent the images rendered from a pseudo view using the current $\mathbf{\Sigma}$-model and $\mathbf{\Delta}$-model, respectively.
+where $\lambda=0.2$ is a predefined weight, $\mathcal{L}_{\text{D-SSIM}}$ denotes a D-SSIM term [21], and $(\mathbf{I}^{t}_{\Sigma},\mathbf{I}^{t}_{\Delta})$ represent the images rendered from a pseudo view using the current $\mathbf{\Sigma}$-model and $\mathbf{\Delta}$-model, respectively.
 We also utilize a co-pruning strategy [47] to enhance the regularization.
 Therefore, the $\mathbf{\Sigma}$-model aggregates information from different samples in the parameter space, serving as an ensemble of 3DGS models.
-The final loss function during training is defined as:
+The final loss function during training is defined as
 $\mathcal{L}=\mathcal{L}_{\text{RGB}}+\gamma\mathcal{L}_{r},$ (14)
 where $\gamma=1$ by default and $\mathcal{L}_{\text{RGB}}$ represents a photometric loss over the training views.
-At testing, we keep the $\mathbf{\Sigma}$-model for novel view synthesis.
-Notably, in contrast to CoR-GS, we derive these samples from a single $\mathbf{\Delta}$-model without additional training costs.
-As a result, our SE-GS can be efficiently scaled up to a large number of 3DGS models, enabling more effective regularization.
-Moreover, our regularization is carried out on pseudo views in a self-supervised manner, which is thus independent of additional ground truth such as monocular depth [23, 53].
+During testing, we keep the $\mathbf{\Sigma}$-model for novel view synthesis.
+Notably, compared with the previous approach [47], our SE-GS inherently encodes a large number of 3DGS models without significant computational overhead, thereby enabling effective regularization efficiently.
+Furthermore, the regularization is carried out on pseudo views in a self-supervised manner, making it independent of additional information such as depth [24, 53].
 […]
@@
 We conduct experiments on four datasets, i.e., LLFF [27], DTU [20], Mip-NeRF360 [3], and MVImgNet [45].
-On LLFF, DTU, and Mip-NeRF360, we follow the experimental setup introduced in CoR-GS [48], using the same training/testing splits and downsampling rates.
-As suggested in CoR-GS, we mask the background when assessing the quality of novel view synthesis on the DTU dataset.
-Since LLFF, DTU, and Mip-NeRF360 offer a limited number of scenarios, we extend our experiment to a large-scale dataset, i.e., MVImgNet.
+On LLFF, DTU, and Mip-NeRF360, we follow the experimental setup introduced in [47], using the same training/testing splits and downsampling rates.
+As suggested in [47], we mask the background when assessing the quality of novel view synthesis on the DTU dataset.
+Since LLFF, DTU, and Mip-NeRF360 offer a limited number of scenarios, we extend our experiments to a large-scale dataset, i.e., MVImgNet.
 We randomly sample 50 scenes from this dataset and resize the longest side of each image to 512.
 The Gaussian parameters for all evaluated methods are initialized based on the point clouds obtained from COLMAP [36] to ensure a fair comparison.
-It is worth noting that COLMAP fails on certain scenes within the DTU dataset in the setting of sparse views, where we instead randomly initialize the point clouds.
+Note that COLMAP fails on certain DTU scenes in the sparse-view setting.
+In these cases, we instead randomly initialize the point cloud.
 Please refer to the supplementary material for more details on the setup.
@@
 We train our SE-GS for 10,000 iterations on the LLFF, DTU, and MVImgNet datasets, and for 30,000 iterations on Mip-NeRF360.
-In the uncertainty-aware perturbing mechanism, we employ 24 image buffers with a buffer size of 3 and perturb the $\mathbf{\Delta}$-model with an interval of 500 iterations.
+In our uncertainty-aware perturbation mechanism, we employ $M=24$ image buffers with a buffer size of $S=3$ and perturb the $\mathbf{\Delta}$-model every 500 iterations.
 The noise $\delta_{t}$ in Eq.
-10 is sampled from a standard normal distribution, i.e., $\delta_{t}\in\mathcal{N}(0,1)$.
+10 is sampled from a normal distribution with a mean of 0, and the standard deviation is adaptively adjusted based on the magnitude of the Gaussian parameters.
 More details are provided in the supplementary material.
@@
 4.2 Quantitative Results
-We list the quantitative results of the evaluated approaches on the LLFF, DTU, Mip-NeRF360, and MVImgNet datasets in Table 1, Table 2, Table 3, and Table 4, respectively.
+We provide the quantitative results of the evaluated approaches on the LLFF, DTU, Mip-NeRF360, and MVImgNet datasets in Table 1, Table 2, Table 3, and Table 4, respectively.
 The best, second-best, and third-best results in each column are highlighted in red, orange, and yellow, respectively.
 For MVSplat [7], we employ the pretrained model released by the authors and evaluate it on DTU.
-Notably, the per-scene optimized 3DGS methods significantly outperform MVSplat, highlighting the advantages of scene-level radiance field modeling for high-quality novel view synthesis.
-The performance of both NeRF and 3DGS methods consistently declines as the number of training views decreases, underscoring the challenge of few-shot novel view synthesis.
+Notably, the per-scene optimized 3DGS methods significantly outperform MVSplat, highlighting the advantage of scene-level radiance field modeling for high-quality novel view synthesis.
+For these scene-level approaches, the performance of both NeRF and 3DGS methods consistently declines as the number of training views decreases, underscoring the challenge of few-shot novel view synthesis.
 In this context, our method surpasses all competitors across the four datasets.
-Note that in the scenario of sparse views, the available information in the training data is limited, making it more challenging to improve performance compared to the dense-view setting.
+Note that in the sparse-view scenario, the available information in the training data is limited, making it more challenging to improve performance compared to the dense-view setting.
 Specifically, on the LLFF and DTU datasets, we train the methods using 3, 6, and 9 views.
 Our method achieves an improvement of $0.34$ in terms of PSNR with 3 views on LLFF.
-The improvement is more significant on DTU, reaching a gain of $0.59$.
-Since the images in the Mip-NeRF360 dataset are captured from diverse viewpoints, we sample more training views, i.e., 12 and 24, following the setting used in CoR-GS.
+The improvement is larger on DTU, reaching a gain of $0.59$.
+Since the images in the Mip-NeRF360 dataset are captured from diverse viewpoints, we sample more training views, i.e., 12 and 24, following the setting used in [47].
 On this dataset, our SE-GS achieves the highest PSNR and SSIM, as well as the lowest LPIPS and AVGE, demonstrating superior NVS quality under the challenges of sparse views with large-scale viewpoint variations.
 Additionally, to further assess the robustness across diverse scenarios, we conduct an experiment on 50 scenes sampled from the MVImgNet dataset.
-As reported in Table 4, our method obtains consistently better results than previous state-of-the-art approaches when trained with 5, 7, and 10 views.
+As reported in Table 4, our method yields consistently better results than previous state-of-the-art approaches when trained with 5, 7, and 10 views.
 Our SE-GS also demonstrates better compatibility with relatively dense views.
 For instance, the PSNR improvement increases from $0.26$ to $0.59$ as the number of training views varies from 5 to 10.
-Methods like DNGaussian [23] and FSGS [53], which leverage auxiliary data terms sometimes outperform 3DGS.
-However, in some cases, such as 9 views on the LLFF dataset, the PSNR of FSGS becomes worse than 3DGS.
+The methods that leverage auxiliary data terms, such as DNGaussian [24] and FSGS [53], sometimes outperform 3DGS.
+However, in some cases, such as with 9 views on LLFF, the PSNR of FSGS becomes worse than 3DGS.
 Since the data acquired through off-the-shelf approaches might be unreliable, the quality of the NVS results is consequently affected.
@@
 However, the improvement is still not consistent, as CoR-GS shows worse LPIPS than 3DGS when trained with 10 views on the MVImgNet dataset.
-In contrast, our SE-GS beats 3DGS in all cases, demonstrating superior stability.
-Our method is capable of avoiding specific local optima since the $\mathbf{\Sigma}$-model aggregates information from diverse temporal samples in the presented self-ensembling mechanism.
-4.3 Qualitative results.
+By contrast, our SE-GS beats 3DGS in all cases, demonstrating superior stability.
+Our method is capable of avoiding specific local optima as the $\mathbf{\Sigma}$-model aggregates information from diverse temporal samples in the presented self-ensembling mechanism.
+4.3 Qualitative Results
 Figure 5: Qualitative results. The methods are trained on sparse views and the renderings of novel views are illustrated. The images are selected from the DTU and MVImgNet datasets.
 Fig.
-5 presents qualitative comparisons among 3DGS, CoR-GS, and our SE-GS, illustrating renderings from novel views.
+5 presents qualitative comparisons of 3DGS, CoR-GS, and our SE-GS, showing renderings from novel views.
 The models are trained on the DTU and MVImgNet datasets with 3 and 5 posed images, respectively.
-Our SE-GS produces fewer visual artifacts compared to the other methods, achieving better robustness against the challenge of few-shot training views.
-Moreover, our method captures finer details, particularly in areas with complex and repeated textures such as the pineapple and watch strap.
-This demonstrates the effectiveness of our self-ensembling strategy in improving both the visual quality and stability of novel view neural rendering under limited-view conditions.
+Our SE-GS produces fewer visual artifacts than the other methods, achieving better robustness against the challenge of few-shot training views.
+Moreover, our method captures finer details, particularly in areas with complex and repeated textures, such as the pineapple and watch strap.
+This demonstrates the effectiveness of our self-ensembling strategy in improving both the visual quality and stability of novel view neural rendering under sparse-view conditions.
 (a) PSNR
@@
 Figure 6: Comparison with CoR-GS.
-CoR-GS [48] is trained with different numbers of Gaussian models.
+CoR-GS [47] is trained with different numbers of 3DGS models.
 The number varies from 2 to 8.
-PSNR on LLFF with 3 training views is utilized as the metric to evaluate the NVS quality.
+The PSNR on LLFF with 3 training views is reported to evaluate the NVS quality.
 Iterations/s refers to the number of training iterations completed per second, indicating the training speed.
 4.4 Analysis
-To shed more light on the effectiveness and efficiency of our SE-GS, we develop a comprehensive analysis of the introduced self-ensembling mechanism.
+To shed more light on the effectiveness and efficiency of our SE-GS, we perform a comprehensive analysis of the introduced self-ensembling mechanism.
 The analysis starts by comparing our approach with CoR-GS trained with varying numbers of Gaussian models.
@@
 6.
-As illustrated in Fig.
+As shown by Fig.
 6(a), the PNSR of CoR-GS increases as more Gaussian models are trained.
 Since adding more Gaussian models enhances the regularization effect in CoR-GS, this observation highlights the potential of improving 3DGS in the sparse-view setting by strengthening the regularization process.
-The PSNR reaches the peak with 6 Gaussian models, indicating an upper bound when using a limited number of models.
-In contrast, our method perturbs the $\mathbf{\Delta}$-model based on the uncertainties computed from dynamically updated image buffers, which results in much more diverse samples of Gaussian models compared to CoR-GS.
+The PSNR reaches a peak with 6 Gaussian models, indicating an upper bound of CoR-GS.
+Our method perturbs the $\mathbf{\Delta}$-model based on the uncertainties computed from dynamically updated image buffers, which results in more diverse samples of Gaussian models compared to CoR-GS.
 This leads to more effective regularization, as reflected by the better PSNR score.
 Moreover, Fig.
-6(b) shows the training speed measured by the number of training iterations completed per second.
-As witnessed in this figure, the speed of CoR-GS significantly drops as the number of Gaussian models increases.
+6(b) shows the training speed measured as the number of training iterations completed per second.
+The speed of CoR-GS significantly drops as the number of Gaussian models increases.
 This limitation poses a challenge in scaling CoR-GS up to a large number of Gaussian models.
-Note that in our pipeline, only two models are trained, i.e., the $\mathbf{\Delta}$-model and $\mathbf{\Sigma}$-model.
+By contrast, in our pipeline, only two models are trained, i.e., the $\mathbf{\Delta}$-model and $\mathbf{\Sigma}$-model.
 The temporal samples are generated through perturbations.
-The result in Fig.
-6(b) shows that we achieve a comparable training speed to CoR-GS with 2 models.
-Only a negligible additional cost is incurred due to the perturbation process.
-It is worth noting that only the $\mathbf{\Sigma}$-model is retained for neural rendering during inference, so the testing speed of SE-GS is the same as that of the original 3DGS.
+We achieve a comparable training speed to CoR-GS with 2 models, showing that the perturbation process incurs a negligible additional cost.
+Note that only the $\mathbf{\Sigma}$-model is retained for NVS during inference, so the testing speed of SE-GS is the same as that of the original 3DGS.
 Consequently, our method is capable of efficiently enhancing the performance of 3DGS with sparse training views.
@@
 Table 5: Effectiveness of the uncertainty-aware perturbations.
-Random means we add random perturbations to all Gaussians in the $\mathbf{\Delta}$-model.
-Gradient-Aware indicates an alternative in which we perturb the Gaussians based on gradients.
+Random indicates that we add random perturbations to all Gaussians in the $\mathbf{\Delta}$-model.
+Gradient-Aware denotes an alternative in which we perturb the Gaussians based on gradients.
 All methods are trained on LLFF with 3 views for 2000 iterations and the PSNR is utilized as the metric.
-Furthermore, we assess the effectiveness of our uncertainty-aware perturbing strategy, comparing it with two alternatives.
-Specifically, as we introduced in Sec.
-3, the straightforward method perturbing the $\mathbf{\Delta}$-model is to add random noise to all Gaussians in the model.
+Finally, we assess the effectiveness of our uncertainty-aware perturbation strategy, comparing it with two alternatives.
+Specifically, as introduced in Sec.
+3, a straightforward method for perturbing the $\mathbf{\Delta}$-model is to add random noise to the parameters of all Gaussians in the model.
 We denote this approach as Random.
-Building upon this baseline, an alternative approach, referred to as Gradient-Aware, identifies unreliable Gaussians by analyzing their gradients.
+Building upon this baseline, an alternative approach, referred to as Gradient-Aware, identifies unreliable Gaussians by analyzing the gradients.
 Specifically, we replace the indicator function in Eq.
-11 with one based on gradient magnitude.
-Gaussians with gradients exceeding a threshold are perturbed.
+11 with one based on gradient magnitudes.
+Gaussians with gradient magnitudes exceeding a threshold are perturbed.
 In practice, we adopt the same threshold as used in density control.
@@
 5 Conclusion
-In this paper, we have introduced a new self-ensembling mechanism that enhances the novel view synthesis of 3DGS with few-shot training images.
-We tackle the challenge of overfitting by training an ensemble model named $\mathbf{\Sigma}$-model with the guidance of diverse temporal samples from the Gaussian parameter space.
-We introduce an effective sampling strategy that perturbs a $\mathbf{\Delta}$-model based on the uncertainties computed from dynamically updated buffers of pseudo-view renderings.
+In this paper, we have presented a new self-ensembling mechanism that enhances the novel view synthesis of 3DGS with sparse training images.
+We have tackled the challenge of overfitting by training an ensemble named $\mathbf{\Sigma}$-model with the guidance of diverse temporal models sampled from the Gaussian parameter space.
+To obtain such samples, we have introduced an effective strategy that perturbs a $\mathbf{\Delta}$-model based on uncertainties computed from dynamically updated buffers of pseudo-view renderings.
 We have conducted experiments on LLFF, DTU, Mip-NeRF360, and MVImgNet, as well as a comprehensive analysis of our approach.
 The experimental results have demonstrated the effectiveness, stability, and efficiency of our SE-GS.
+In future work, we plan to incorporate an outlier identifier into our perturbation mechanism to reduce the impact of unreliable samples in regularization and facilitate the self-ensembling process.
 Acknowledgment.
 This work was funded in part by the Swiss National Science Foundation via the Sinergia grant CRSII5-180359 and the Swiss Innovation Agency (Innosuisse) via the BRIDGE Discovery grant 40B2-0_194729.
+[…]
+Supplementary Material
+Details on Uncertainty-Aware Perturbation
+Recall that in Eq.
+11 of the main paper, we associate the reliability of Gaussians with the pixel-level uncertainty scores.
+In practice, for each uncertainty map $\hat{\mathbf{U}}$, we define $\tau$ as
+$\tau$ $=\text{max}(\hat{\mathbf{U}}_{\text{sorted}}\left[\left\lceil r*|\hat{\mathbf{U}}_{\text{sorted}}|\right\rceil\right],\theta),$ (15)
+[…]
+In Eq.
+10, we sample the noise $\delta_{t}$ from a normal distribution.
+As an example, we elaborate on the perturbation process for the 3D position, which is similar for the other Gaussian parameters.
+In this context, we define $\delta_{t}$ as $\delta_{t}\in\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ where $\mathbf{I}\in\mathbb{R}^{3\times 3}$ is an identity matrix.
+In practice, we adaptively adjust $\sigma$ based on the magnitude of the parameters, as formulated by
+$\sigma=\omega\frac{1}{N}\sum_{i=1}^{N}{||\mathbf{\mu}_{i}||}_{1},$ (16)
+where $\omega$ is a scalar that controls the noise level and ${||\cdot||}_{1}$ denotes the L1 norm of the Gaussian parameter.
+The value of $\omega$ decays from 0.08 to 0.02, following a decay function introduced in [12].
+Setup on DTU
+As we mentioned in the main paper, COLMAP fails in some scenes on DTU when using sparse-view images.
+In our experiments, we randomly initialize the point cloud in these scenes.
+Specifically, we use random initialization for scan8, scan40, and scan110 with 3 training views, and for scan21 with 6 training views.
+Ablation Studies
+To shed more light on the impact of noise during the perturbation, we conduct comprehensive ablation studies on the perturbation interval and noise level.
+Figure 7: Ablation study on the perturbation interval. PSNR values on testing views throughout training are reported. The interval varies from 100 to 900.
+Figure 8: Ablation study on the noise level. We evaluate our method using different values of $\omega$, each corresponding to a distinct noise level.
+In our experiments, we perturb the $\mathbf{\Delta}$-model every 500 iterations by default.
+To evaluate the effect of different intervals, we test our method on the LLFF dataset with 3 training views, using intervals ranging from 100 to 900.
+As shown in Fig.
+7, the curves represent the PSNR values obtained on testing views at different training iterations.
+Perturbing the $\mathbf{\Delta}$-model too frequently, e.g., with an interval of 100, significantly degrades performance, indicating a negative impact on our method.
+Conversely, large intervals, such as 900, reduce the number of temporal samples, thereby weakening the self-ensembling effect and leading to decreased performance compared to the default setting.
+Moreover, we analyze the effect of the noise level by conducting experiments with varying values of $\sigma$ in Eq.
+16.
+Specifically, we adjust the value of $\omega$ from 0.01 to 0.3, corresponding to increasing noise levels.
+We train our method with each specific noise level on the LLFF dataset using 3 training views.
+The results on testing views at different training iterations are shown in Fig.
+8.
+Strong perturbations, such as those with $\omega$ values of 0.3 and 0.2, result in significantly worse performance compared to other settings.
+In these cases, the temporal models are sampled too far from the $\mathbf{\Delta}$-model in the Gaussian parameter space, reducing the consistency among temporal samples.
+The performance is also limited when a small $\omega$, such as 0.01, is used.
+In this scenario, all temporal models closely resemble the $\mathbf{\Delta}$-model, diminishing the benefits of self-ensembling.
+In contrast to using a fixed $\omega$, we adopt a decay function, where $\omega$ dynamically varies from 0.08 to 0.02 during training.
+This strategy achieves a good trade-off between the consistency and diversity of the temporal samples.
+Figure 9: Number of perturbed Gaussians throughout training. The curve illustrates the percentage of perturbed Gaussians in the $\mathbf{\Delta}$-model at different training iterations.
+To better understand the presented perturbation mechanism, in Fig.
+9, we illustrate the percentage of perturbed Gaussians in the $\mathbf{\Delta}$-model.
+Note that the threshold $\tau$ in Eq.
+15 is defined adaptively.
+Therefore, the percentage dynamically varies throughout training.
+This dynamic behavior balances the diversity and reliability of the temporal models, thereby enhancing the self-ensembling process.
+More Visualization Results
+We provide more visualizations including pseudo-view renderings of the $\mathbf{\Sigma}$-model and the $\mathbf{\Delta}$-model, uncertainty maps, and NVS results.
+For more details, please refer to the uploaded videos.
+Notably, the videos of pseudo-view renderings and uncertainty maps showcase the results across different training iterations.

### v3

@@
 Abstract
-3D Gaussian Splatting (3DGS) has demonstrated remarkable effectiveness for novel view synthesis (NVS).
-However, the 3DGS model tends to overfit when trained with sparse posed views, limiting its generalization ability to novel views.
-In this paper, we alleviate the overfitting problem, presenting a Self-Ensembling Gaussian Splatting (SE-GS) approach.
-Our method encompasses a $\mathbf{\Sigma}$-model and a $\mathbf{\Delta}$-model.
-The $\mathbf{\Sigma}$-model serves as an ensemble of 3DGS models that generates novel-view images during inference.
-We achieve the self-ensembling by introducing an uncertainty-aware perturbation strategy at the training state.
-We complement the $\mathbf{\Sigma}$-model with the $\mathbf{\Delta}$-model, which is dynamically perturbed based on the uncertainties of novel-view renderings across different training steps.
-The perturbation yields diverse temporal samples in the Gaussian parameter space without additional training costs.
-The geometry of the $\mathbf{\Sigma}$-model is regularized by penalizing discrepancies between the $\mathbf{\Sigma}$-model and these temporal samples.
-Therefore, our SE-GS conducts an effective and efficient regularization across a large number of 3DGS models, resulting in a robust ensemble, the $\mathbf{\Sigma}$-model.
-Our experimental results on the LLFF, Mip-NeRF360, DTU, and MVImgNet datasets show that our approach improves NVS quality with few-shot training views, outperforming existing state-of-the-art methods.
+3D Gaussian Splatting (3DGS) has demonstrated remarkable effectiveness in novel view synthesis (NVS).
+However, 3DGS tends to overfit when trained with sparse views, limiting its generalization to novel viewpoints.
+In this paper, we address this overfitting issue by introducing Self-Ensembling Gaussian Splatting (SE-GS).
+We achieve self-ensembling by incorporating an uncertainty-aware perturbation strategy during training.
+A $\mathbf{\Delta}$-model and a $\mathbf{\Sigma}$-model are jointly trained on the available images.
+The $\mathbf{\Delta}$-model is dynamically perturbed based on rendering uncertainty across training steps, generating diverse perturbed models with negligible computational overhead.
+Discrepancies between the $\mathbf{\Sigma}$-model and these perturbed models are minimized throughout training, forming a robust ensemble of 3DGS models.
+This ensemble, represented by the $\mathbf{\Sigma}$-model, is then used to generate novel-view images during inference.
+Experimental results on the LLFF, Mip-NeRF360, DTU, and MVImgNet datasets demonstrate that our approach enhances NVS quality under few-shot training conditions, outperforming existing state-of-the-art methods.
 The code is released at: project page.
@@
 To illustrate this issue, in Fig.
-2, we evaluate the 3DGS model trained on sparse images with different numbers of iterations.
+2, we evaluate 3DGS trained on sparse images with different numbers of iterations.
 The performance on the training data consistently improves as the number of iterations increases, while the testing results deteriorate after 2000 iterations.
@@
 Figure 2: Overfitting in 3D Gaussian Splatting with sparse training views.
-2(a) and 2(b) illustrate the performance of 3DGS on training and testing views, respectively.
-Each curve represents the PSNR values across different training iterations.
-To mitigate overfitting, self-ensembling [11, 22, 29] has been highlighted as an effective strategy in tasks such as detection [48] and segmentation [25].
+(a) and (b) illustrate the performance of 3DGS on training and testing views, respectively.
+Each curve represents the PSNR values across training iterations.
+To mitigate overfitting, ensembling [23, 22, 10] has been highlighted as an effective strategy in detection [49] and segmentation [26].
 Nevertheless, its use for NVS remains unstudied, and how to exploit it in a 3DGS formalism is an open question.
 Therefore, in this paper, we bridge this gap, introducing a new 3DGS method that enhances the quality of novel view synthesis with sparse training views via self-ensembling.
-Our method consists of two models, a $\mathbf{\Sigma}$-model and a $\mathbf{\Delta}$-model.
-The $\mathbf{\Sigma}$-model constitutes an ensemble of 3DGS models, which is used to generate novel views during inference.
-A straightforward way to achieve ensembling would be to utilize multiple 3DGS models that correspond to different samples in the Gaussian parameter space, thereby increasing the likelihood of reaching a more robust solution.
-Nevertheless, as we will demonstrate in Sec.
-4, the computational cost of training multiple 3DGS models limits the ability to produce sufficiently diverse samples.
-By contrast, we complement the $\mathbf{\Sigma}$-model with a single $\mathbf{\Delta}$-model and obtain diverse samples by perturbing the $\mathbf{\Delta}$-model.
-Specifically, instead of applying random perturbation, which often results in invalid samples, we perturb the $\mathbf{\Delta}$-model based on uncertainties derived from the training data.
-We store images generated by rendering the $\mathbf{\Delta}$-model from a set of pseudo views at different training iterations in buffers and calculate pixel-level uncertainties across images within each buffer.
-We then add random noise to the Gaussian parameters of the $\mathbf{\Delta}$-model associated with pixels that have high uncertainty scores.
-Consequently, our perturbation strategy generates uncertainty-aware samples without introducing significant training overhead.
-Building upon the uncertainty-aware samples derived from the $\mathbf{\Delta}$-model, we construct an ensemble by leveraging a self-ensembling mechanism.
-We train the $\mathbf{\Sigma}$-model over the training views without the aforementioned perturbation and regularize the geometry of the $\mathbf{\Sigma}$-model by minimizing its discrepancies with the perturbed $\mathbf{\Delta}$-model.
-These discrepancies are measured via a photometric loss between images synthesized from the pseudo views.
-Since the regularization is performed based on diverse samples in the Gaussian parameter space, the self-ensembling process enhances the robustness of the $\mathbf{\Sigma}$-model, thereby improving its generalization ability to novel views.
-Moreover, our self-ensembling technique is independent of any additional ground-truth signals as the regularization is carried out in a self-supervised manner.
-We conduct experiments on multiple datasets, including LLFF [27], DTU [20], Mip-NeRF360 [3], and MVImgNet [45], with limited numbers of training views.
+A straightforward way to achieve ensembling would be to train multiple 3DGS models and combine the corresponding predictions as the final result during testing.
+However, as we will demonstrate in Sec.
+4, this method is computationally expensive, and the resulting models lack sufficient diversity to ensure effective ensembling.
+In contrast, we present an effective and efficient self-ensembling mechanism.
+In ensemble learning [11, 30], self-ensembling is typically achieved by introducing perturbations during training through augmentation or dropout.
+Inspired by its success, we introduce an uncertainty-aware perturbation strategy for 3DGS.
+Specifically, we train a $\mathbf{\Delta}$-model on available RGB images.
+During training, we perturb the $\mathbf{\Delta}$-model based on uncertainties derived from the renderings.
+To compute these uncertainties, we store images rendered from the $\mathbf{\Delta}$-model at pseudo views across different training iterations in buffers and calculate pixel-level uncertainties within each buffer.
+A perturbed model is then obtained by adding random noise to the Gaussian parameters of the $\mathbf{\Delta}$-model associated with pixels that have high uncertainty scores.
+Since all perturbed models are derived from the $\mathbf{\Delta}$-model rather than being trained from scratch, our perturbation strategy generates diverse models without incurring significant computational overhead.
+In addition to the $\mathbf{\Delta}$-model, we train a $\mathbf{\Sigma}$-model on the training views without the perturbation.
+The self-ensembling is achieved by minimizing the discrepancies between the $\mathbf{\Sigma}$-model and perturbed models.
+These discrepancies are measured via a photometric loss between images synthesized at the pseudo views.
+This self-ensembling is performed based on diverse models, thereby improving the robustness and generalization of the resulting ensemble, the $\mathbf{\Sigma}$-model.
+Moreover, our self-ensembling is independent of any additional ground-truth signals as it is carried out in a self-supervised manner.
+During testing, the $\mathbf{\Sigma}$-model is employed for novel view synthesis.
+We conduct experiments on multiple datasets, including LLFF [28], DTU [20], Mip-NeRF360 [3], and MVImgNet [46], with sparse training views.
 Our SE-GS achieves the best performance across all datasets in the sparse-view setting, surpassing the state-of-the-art approaches.
@@
 Recently, Neural Radiance Field (NeRF) [29, 42, 33, 3] has brought groundbreaking advancements to novel view synthesis.
-NeRF learns the density and color information of a scene or object from multi-view posed images, enabling the synthesis of realistic images from novel views via volume rendering [9].
-However, most NeRF-based methods are inefficient [18] during training and inference because they require numerous MLP queries in the rendering process.
+However, most NeRF-based methods are inefficient [18] because they require numerous MLP queries in the rendering process.
 This limitation makes it challenging to support tasks with real-time requirements.
@@
 3DGS relies on explicit representations [36] that allow for faster rendering and training [9].
-While initially introduced as a scene-specific model, some 3DGS variants [40, 5, 26, 7] explore the generalization ability of 3DGS towards novel scenes, optimizing the Gaussian parameters based on learnable cost volumes [15].
+While initially introduced as a scene-specific model, some 3DGS variants [41, 5, 27, 7] explore the generalization of 3DGS towards novel scenes, optimizing the Gaussian parameters based on learnable cost volumes [15].
 In this paper, we stick to the vanilla 3DGS rendering pipeline, focusing on the scene-level radiance field modeling.
@@
 However, training additional 3DGS models incurs a significant computational cost, making it impractical to scale up to a large number of 3DGS models for stronger regularization.
-Ensembling.
+Ensemble learning.
 Ensembling has been evidenced as a powerful technique in machine learning [53, 1, 14] to improve model robustness and generalization by aggregating predictions from multiple models.
@@
 Figure 3: Pipeline of the presented SE-GS.
-We tackle the overfitting problem in sparse-view scenarios by incorporating a self-ensembling mechanism into the 3DGS framework.
-Our self-ensembling builds upon diverse samples in the Gaussian parameter space.
-These samples are derived from the $\mathbf{\Delta}$-model via an uncertainty-aware perturbation strategy.
-Specifically, we store images rendered from pseudo views at different training steps in buffers, from which we compute pixel-level uncertainties.
-We then perturb the Gaussians overlapping the pixels with high uncertainties, as highlighted as red ellipses.
-Self-ensembling over the resulting samples is achieved by training a $\mathbf{\Sigma}$-model with a regularization that penalizes the discrepancies of the $\mathbf{\Sigma}$-model and the perturbed $\mathbf{\Delta}$-model.
-During inference, novel view synthesis is performed using the $\mathbf{\Sigma}$-model.
-3.1 Preliminaries and Motivations
+We tackle the overfitting problem in sparse-view scenarios by incorporating a self-ensembling mechanism into 3DGS.
+We jointly train a $\mathbf{\Delta}$-model and a $\mathbf{\Sigma}$-model.
+During training, we store pseudo-view renderings of the $\mathbf{\Delta}$-model in buffers, from which we compute pixel-level uncertainties.
+The Gaussians of the $\mathbf{\Delta}$-model overlapping the pixels with high uncertainties are perturbed, as highlighted as red ellipses, which leads to a perturbed model.
+We then achieve self-ensembling by penalizing the discrepancies between the $\mathbf{\Sigma}$-model and the perturbed models.
+During inference, the resulting ensemble, the $\mathbf{\Sigma}$-model, is used for novel view synthesis.
+3.1 Preliminaries
 3D Gaussian Splatting [21] represents a scene as a collection of Gaussians.
@@
 $\mathcal{N}_{3d}^{i}(\mathbf{x})=e^{-\frac{1}{2}(\mathbf{x}-\mathbf{\mu}_{i})^{T}\Sigma_{i}^{-1}(\mathbf{x}-\mathbf{\mu}_{i})}.$ (1)
-The parameters of the Gaussians are then learned from the input views.
 To ensure that $\Sigma_{i}$ remains positive semi-definite throughout optimization, it is decomposed into two learnable components as $\Sigma_{i}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T}$, where $\mathbf{R}$ is a rotation matrix and $\mathbf{S}$ stands for a scaling matrix.
@@
 where $M$ is the number of Gaussians covering the pixel, $\mathbf{c}_{i}$ denotes the color of the Gaussian derived from the SH coefficients, and $\alpha_{i}$ is computed from the 2D covariance matrices and opacity scores.
-For better convergence, the Gaussian parameters are initialized based on 3D points obtained using structure-from-motion (SFM) techniques [35, 36].
-These parameters are optimized via gradient descent with a photometric loss [21] where the posed training images serve as ground truth.
-In this paper, we focus on sparse-view scenarios, where only few-shot training images are provided.
-We denote the 3DGS model trained on these images as $\mathcal{G}$, which represents a sample in the Gaussian parameter space.
+For better convergence, the Gaussian parameters are initialized based on 3D points obtained using structure-from-motion techniques [36, 37].
+These parameters are optimized with a photometric loss [21] where the posed training images serve as ground truth.
+3.2 Motivations
+In this paper, we focus on sparse-view scenarios, where few-shot images are provided.
+We denote the 3DGS model trained on these images as $\mathcal{G}$.
 In this setting, $\mathcal{G}$ is prone to overfitting the training data and thereby getting trapped in a suboptimal solution, which ultimately degrades the quality of the synthesized novel views.
-To handle the overfitting problem, a promising solution is to train a set of 3DGS models $\{\mathcal{G}_{1},\mathcal{G}_{2},...,\mathcal{G}_{k}\}$ that correspond to different samples.
-By applying regularization across this set, each model is guided away from specific local optima, facilitating a more thorough exploration of the parameter space.
+A promising approach to mitigate overfitting is ensemble learning [13], which has been utilized to enhance robustness and generalization in the literature.
+Rather than depending on a single model, ensemble learning aggregates predictions from multiple models, thereby stabilizing the final output.
+To achieve this, a straightforward method is to jointly train a set of 3DGS models $\{\mathcal{G}_{1},\mathcal{G}_{2},...,\mathcal{G}_{k}\}$.
 As we will report in Sec.
-4, the performance on testing data improves as $k$ grows.
-However, training multiple 3DGS models incurs a significant computational cost, which becomes unaffordable with a large $k$.
-This limitation prevents the method from scaling up for better performance.
-Consequently, we propose a novel training strategy based on self-ensembling [11, 22], which enhances 3DGS for few-shot NVS with negligible additional training overhead.
-The pipeline is illustrated in Fig.
-3 and the details will be elaborated in the remainder of this section.
+4, by aggregating their predictions, the final results are more robust, and the NVS performance improves as $k$ grows.
+However, training multiple 3DGS models incurs significant computational costs, making it impractical for large $k$.
+Moreover, our experimental findings indicate that the trained 3DGS models are not diverse enough to support effective ensembling.
+Consequently, we propose a novel self-ensembling paradigm based on an uncertainty-aware perturbation strategy, enhancing 3DGS for few-shot NVS with negligible additional training overhead.
+The pipeline is shown in Fig.
+3 and the details will be elaborated in this section.
 3.3 Uncertainty-Aware Perturbation
 Figure 4: Buffer update during training.
-For each sampled pseudo view, we dynamically update the buffer that stores the images rendered at different training steps.
+For each sampled pseudo view, we dynamically update the buffer storing the images rendered at different training steps.
 For instance, at training step $t_{T}$, the oldest image $\mathbf{I}_{T-S}$ in the buffer is popped, and the new image $\mathbf{I}_{T}$ is pushed into the buffer.
-An uncertainty map $\mathbf{U}^{t_{T}}$ is computed based on the current buffer, which is then employed to compute a perturbation that results in a new 3DGS model.
-In contrast to generating multiple samples by training separate 3DGS models, we dynamically create temporal samples from a single $\mathbf{\Delta}$-model during training.
-Note that, in this paper, temporal refers to the dynamic nature of samples that vary across training iterations.
+An uncertainty map $\mathbf{U}^{t_{T}}$ is computed based on the current buffer, which is then employed to determine perturbation that results in a new 3DGS model.
+In contrast to training separate 3DGS models from scratch, we dynamically generate diverse models from a single $\mathbf{\Delta}$-model during training.
 Specifically, we train the $\mathbf{\Delta}$-model on the available posed images, following the same optimization and density control strategies introduced in 3DGS [21].
 At each training iteration, the current $\mathbf{\Delta}$-model represents a specific sample in the Gaussian parameter space based on the information contained in the training data.
-A temporal 3DGS model is then sampled from the neighboring region of the current $\mathbf{\Delta}$-model in the parameter space by perturbing the $\mathbf{\Delta}$-model as
+A new 3DGS model is then created by perturbing the $\mathbf{\Delta}$-model as
 $\hat{G}_{\Delta}^{t}=G_{\Delta}^{t}+\delta_{t},$ (4)
-where $G_{\Delta}^{t}$ stands for a Gaussian in the $\mathbf{\Delta}$-model at training step $t$, $\delta_{t}$ indicates the perturbation, and $\hat{G}_{\Delta}^{t}$ denotes a perturbed Gaussian in the temporal model.
-To achieve the perturbation, a naive approach is to sample random noise from a normal distribution, i.e., $\delta_{t}\in\mathcal{N}(\mu_{t},\sigma_{t}^{2})$.
-However, such perturbation is independent of the training data, and this method thus tends to yield invalid samples.
+where $G_{\Delta}^{t}$ stands for a Gaussian in the $\mathbf{\Delta}$-model at training step $t$, $\delta_{t}\in\mathcal{N}(\mathbf{\mu_{t}},\mathbf{\sigma}^{2}_{t})$ indicates the noise, and $\hat{G}_{\Delta}^{t}$ denotes a perturbed Gaussian in the perturbed model.
+This naive approach adds random noise to all Gaussians in the $\mathbf{\Delta}$-model.
+However, as we will demonstrate in Sec.
+4, this method shifts perturbed models too far from the $\mathbf{\Delta}$-model, leading to instability and unreliable supervision.
 Therefore, we present an uncertainty-aware perturbation strategy, leveraging the statistics of renderings.
@@
 $\hat{\mathbf{T}}=-\hat{\mathbf{R}}\hat{\mathbf{c}},$ (7)
-where $(\mathbf{R}_{1},\mathbf{R}_{2})$ represent the rotations of the sampled cameras, $(\mathbf{c}_{1},\mathbf{c}_{2})$ indicate the sampled camera centers, SLERP denotes the spherical linear interpolation [4], $\beta$ is a scalar that controls the interpolation, and $(\hat{\mathbf{R}},\hat{\mathbf{T}})$ stand for the camera parameters of the pseudo view.
+where $(\mathbf{R}_{1},\mathbf{R}_{2})$ represent the rotations of the sampled cameras, $(\mathbf{c}_{1},\mathbf{c}_{2})$ indicate the sampled camera centers, SLERP denotes the spherical linear interpolation [4], $\beta$ is a randomly sampled scalar that controls the interpolation, and $(\hat{\mathbf{R}},\hat{\mathbf{T}})$ stand for the camera parameters of the pseudo view.
 As illustrated in Fig.
@@
 $\mathbf{U}=\sqrt{\frac{1}{S}\sum_{i=1}^{S}(\mathbf{I}_{i}-\bar{\mathbf{I}})^{2}},$ (8)
-of the same resolution as the rendered image, i.e., $\mathbf{U}\in\mathbb{R}^{H\times W}$, where $\mathbf{I}_{i}$ represents an image in the buffer, $\bar{\mathbf{I}}$ indicates the mean of these images, and $S$ is the buffer size.
+where $\mathbf{I}_{i}$ represents an image in the buffer, $\bar{\mathbf{I}}$ indicates the mean of these images, and $S$ is the buffer size.
 This uncertainty estimation is carried out over all $M$ sampled pseudo views in parallel.
-In practice, to enhance robustness, we perform local smoothing over each uncertainty map, which yields
-$\hat{\mathbf{U}}(i,j)=\frac{1}{k^{2}}\sum_{m,n}\mathbf{U}(i+m,j+n),$ (9)
-where $k=5$ is the kernel size. Subsequently, we apply an uncertainty-aware perturbation to each of the Gaussians in the $\mathbf{\Delta}$-model as
+To enhance robustness, we perform local smoothing over each uncertainty map, yielding
+$\hat{\mathbf{U}}(i,j)=\frac{1}{k^{2}}\sum_{m,n}\mathbf{U}(m,n),$ (9)
+where $k=5$ is the size of a kernel centered at $(i,j)$ and $(m,n)$ indicate the coordinates of a pixel within the kernel.
+Subsequently, we apply uncertainty-aware perturbation to each Gaussian in the $\mathbf{\Delta}$-model as
 $\hat{G}_{\Delta}^{t}=G_{\Delta}^{t}+\delta_{t}h(G_{\Delta}^{t},\hat{\mathcal{U}}^{t}),$ (10)
-where $\hat{\mathcal{U}}^{t}$ is the set of the uncertainty maps computed from the buffers at the current training step, and $h(\cdot,\cdot)$ is an indicator function defined as
+where $\hat{\mathcal{U}}^{t}=\{\hat{\mathbf{U}}^{t}_{1},\hat{\mathbf{U}}^{t}_{2},\cdots,\hat{\mathbf{U}}^{t}_{M}\}$ is a set of uncertainty maps computed from the buffers at the current training step, and $h(\cdot,\cdot)$ is an indicator function defined as
 $h(G_{\Delta}^{t},\hat{\mathcal{U}}^{t})=\begin{cases}1&\text{if }\max\limits_{(i,j)\in\mathcal{P}(G_{\Delta}^{t})}u_{ij}\geq\tau\\ 0&\text{if }\max\limits_{(i,j)\in\mathcal{P}(G_{\Delta}^{t})}u_{ij}<\tau,\end{cases}$ (11)
-with $\tau$ a predefined threshold, and $u_{ij}$ the uncertainty score of a pixel sampled from the set of pixels $\mathcal{P}$ overlapping with the 2D splats of $G_{\Delta}^{t}$.
+with $\tau$ a predefined threshold, and $u_{ij}$ the uncertainty score of a pixel in a set of pixels $\mathcal{P}$ overlapping with the 2D splats of $G_{\Delta}^{t}$.
 Please refer to the supplementary material for details on defining $\tau$ and identifying $\mathcal{P}$.
-In short, the reliability of each Gaussian is connected with the uncertainties of the renderings, and only the unreliable Gaussians, characterized by high uncertainty scores, are perturbed.
-In practice, we add the perturbation to the 3D positions, 3D rotations, scales, and opacities of the $\mathbf{\Delta}$-model.
-In particular, since 3D rotation is not continuous when expressed as a 3D matrix, we perturb the 6D continuous representation [51] of the rotation, which yields
+In short, the reliability of each Gaussian is connected with the rendering uncertainty, and only the unreliable Gaussians, characterized by high uncertainty scores, are perturbed.
+In practice, we add random noise to 3D positions, 3D rotations, scales, and opacities of the $\mathbf{\Delta}$-model.
+Particularly, since rotation is not continuous when expressed as a 3D matrix, we perturb the 6D continuous representation [52], which is denoted as
 $\hat{\mathbf{R}}^{t}_{\Delta}=f^{-1}(f(\mathbf{R}_{\Delta}^{t})+\delta^{R}_{t}h(G_{\Delta}^{t},\hat{\mathcal{U}}^{t})),\ \ \delta^{R}_{t}\in\mathbb{R}^{6},$ (12)
-where $f(\cdot)$ denotes the mapping from rotation matrix to 6D representation.
+where $f(\cdot)$ indicates the mapping from rotation matrix to 6D representation.
 Fig.
 4 illustrates the dynamic updates of the buffer throughout the training process.
-As the buffer is updated, the uncertainty map varies accordingly, resulting in diverse temporal samples from the $\mathbf{\Delta}$-model.
-It is worth noting that the number of Gaussians in the $\mathbf{\Delta}$-model varies during training due to density control, making it challenging to compute uncertainties directly in the Gaussian parameter space across different training steps.
+As the buffer is updated, the uncertainty map varies accordingly, resulting in diverse perturbed models from the $\mathbf{\Delta}$-model.
+It is worth noting that the number of Gaussians in the $\mathbf{\Delta}$-model varies during training due to density control, making it challenging to assess uncertainties directly in the Gaussian parameter space for newly generated Gaussians.
 Our approach naturally handles this challenge as the statistics on 2D renderings provide a consistent ground to measure uncertainties, regardless of the varying number of Gaussians.
 3.4 Self-Ensembling in 3DGS
-Given the temporal samples around the $\mathbf{\Delta}$-model, the next step in our pipeline is to construct an ensemble from these samples.
-Concretely, we train a separate 3DGS model, named the $\mathbf{\Sigma}$-model, on the posed images without perturbations.
-To prevent the $\mathbf{\Sigma}$-model from being trapped at a suboptimal solution, we guide the training of the $\mathbf{\Sigma}$-model with a regularization [47] formulated as
+Given the perturbed models derived from the $\mathbf{\Delta}$-model, the next step in our pipeline is to construct an ensemble from these models.
+To ensure efficiency, we perform self-ensembling during training.
+Concretely, we train a separate 3DGS model, named the $\mathbf{\Sigma}$-model, on the training views without perturbation.
+Its training is guided with an additional regularization formulated as
 $\mathcal{L}_{r}=(1-\lambda){||\mathbf{I}^{t}_{\Sigma}-\mathbf{I}^{t}_{\Delta}||}_{1}+\lambda\mathcal{L}_{\text{D-SSIM}}(\mathbf{I}^{t}_{\Sigma},\mathbf{I}^{t}_{\Delta}),$ (13)
-where $\lambda=0.2$ is a predefined weight, $\mathcal{L}_{\text{D-SSIM}}$ denotes a D-SSIM term [21], and $(\mathbf{I}^{t}_{\Sigma},\mathbf{I}^{t}_{\Delta})$ represent the images rendered from a pseudo view using the current $\mathbf{\Sigma}$-model and $\mathbf{\Delta}$-model, respectively.
+where $\lambda=0.2$ is a predefined weight, $\mathcal{L}_{\text{D-SSIM}}$ denotes a D-SSIM term [21], and $(\mathbf{I}^{t}_{\Sigma},\mathbf{I}^{t}_{\Delta})$ represent the images rendered from a pseudo view using the current $\mathbf{\Sigma}$-model and the perturbed model, respectively.
 We also utilize a co-pruning strategy [48] to enhance the regularization.
-Therefore, the $\mathbf{\Sigma}$-model aggregates information from different samples in the parameter space, serving as an ensemble of 3DGS models.
+Therefore, the $\mathbf{\Sigma}$-model aggregates information from diverse perturbed models, functioning as an ensemble of 3DGS models.
 The final loss function during training is defined as
@@
 During testing, we keep the $\mathbf{\Sigma}$-model for novel view synthesis.
-Notably, compared with the previous approach [47], our SE-GS inherently encodes a large number of 3DGS models without significant computational overhead, thereby enabling effective regularization efficiently.
-Furthermore, the regularization is carried out on pseudo views in a self-supervised manner, making it independent of additional information such as depth [24, 53].
+Notably, compared with the previous approach [48], our SE-GS inherently encodes diverse 3DGS models without significant computational overhead, thereby enabling efficient and effective regularization.
+Moreover, the regularization is applied on pseudo views in a self-supervised manner, making it independent of additional information such as depth [25, 54].
 […]
-3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view
-Mip-NeRF [2] 16.11 22.91 24.88 0.401 0.756 0.826 0.460 0.213 0.160 0.215 0.090 0.066
-DietNeRF [19] 14.94 21.75 24.28 0.370 0.717 0.801 0.496 0.248 0.183 0.240 0.105 0.073
-RegNeRF [30] 19.08 23.10 24.86 0.587 0.760 0.820 0.336 0.206 0.161 0.149 0.086 0.067
-FreeNeRF [44] 19.63 23.73 25.13 0.612 0.779 0.827 0.308 0.195 0.160 0.134 0.075 0.064
-SparseNeRF [38] 19.86 - - 0.624 - - 0.328 - - 0.127 - -
-3DGS [21] 19.22 23.80 25.44 0.649 0.814 0.860 0.229 0.125 0.096 0.120 0.066 0.051
-DNGaussian [24] 19.12 - - 0.591 - - 0.294 - - 0.132 - -
-FSGS [53] 20.43 24.09 25.31 0.682 0.823 0.860 0.248 0.145 0.122 0.104 0.066 0.054
-CoR-GS [47] 20.45 24.49 26.06 0.712 0.837 0.874 0.196 0.115 0.089 0.101 0.060 0.046
-SE-GS 20.79 24.78 26.36 0.724 0.839 0.878 0.183 0.110 0.084 0.100 0.059 0.045
+3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view
+Mip-NeRF [2] 16.11 22.91 24.88 0.401 0.756 0.826 0.460 0.213 0.160
+DietNeRF [19] 14.94 21.75 24.28 0.370 0.717 0.801 0.496 0.248 0.183
+RegNeRF [31] 19.08 23.10 24.86 0.587 0.760 0.820 0.336 0.206 0.161
+FreeNeRF [45] 19.63 23.73 25.13 0.612 0.779 0.827 0.308 0.195 0.160
+SparseNeRF [39] 19.86 - - 0.624 - - 0.328 - -
+3DGS [21] 19.22 23.80 25.44 0.649 0.814 0.860 0.229 0.125 0.096
+DNGaussian [25] 19.12 22.01 22.62 0.591 0.717 0.741 0.294 0.246 0.244
+FSGS [54] 20.43 24.09 25.31 0.682 0.823 0.860 0.248 0.145 0.122
+CoR-GS [48] 20.45 24.49 26.06 0.712 0.837 0.874 0.196 0.115 0.089
+SE-GS 20.79 24.78 26.36 0.724 0.839 0.878 0.183 0.110 0.084
 Table 1: Results on LLFF with 3, 6, and 9 training views. We highlight the best, second-best, and third-best results in red, orange, and yellow, respectively.
 […]
-3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view
-3DGS [21] 17.67 23.69 26.80 0.804 0.894 0.941 0.158 0.086 0.050 0.112 0.058 0.031
-MVSplat [7] 17.33 16.34 16.22 0.598 0.596 0.587 0.279 0.296 0.304 - - -
-DNGaussian [24] 18.57 - - 0.776 - - 0.178 - - 0.110 - -
-FSGS [53] 17.84 23.68 26.17 0.822 0.905 0.941 0.161 0.096 0.064 0.110 0.056 0.036
-CoR-GS [47] 18.65 24.39 27.38 0.835 0.910 0.950 0.140 0.074 0.045 0.099 0.049 0.028
-SE-GS 19.24 25.28 28.08 0.857 0.924 0.958 0.132 0.073 0.043 0.089 0.045 0.026
+3-view 6-view 9-view 3-view 6-view 9-view 3-view 6-view 9-view
+3DGS [21] 17.67 23.69 26.80 0.804 0.894 0.941 0.158 0.086 0.050
+MVSplat [7] 17.33 16.34 16.22 0.598 0.596 0.587 0.279 0.296 0.304
+DNGaussian [25] 18.57 22.56 25.25 0.776 0.862 0.917 0.178 0.114 0.077
+FSGS [54] 17.84 23.68 26.17 0.822 0.905 0.941 0.161 0.096 0.064
+CoR-GS [48] 18.65 24.39 27.38 0.835 0.910 0.950 0.140 0.074 0.045
+SE-GS 19.24 25.28 28.08 0.857 0.924 0.958 0.132 0.073 0.043
 Table 2: Results on DTU with 3, 6, and 9 training views.
@@
 We conduct experiments on four datasets, i.e., LLFF [28], DTU [20], Mip-NeRF360 [3], and MVImgNet [46].
-On LLFF, DTU, and Mip-NeRF360, we follow the experimental setup introduced in [47], using the same training/testing splits and downsampling rates.
+On LLFF, DTU, and Mip-NeRF360, we follow the experimental setup introduced in [48], using the same training/testing splits.
 As suggested in [48], we mask the background when assessing the quality of novel view synthesis on the DTU dataset.
@@
 We randomly sample 50 scenes from this dataset and resize the longest side of each image to 512.
-The Gaussian parameters for all evaluated methods are initialized based on the point clouds obtained from COLMAP [36] to ensure a fair comparison.
-Note that COLMAP fails on certain DTU scenes in the sparse-view setting.
+Notably, in our experiments, the Gaussian parameters for all evaluated methods are initialized based on the same point clouds obtained from COLMAP [37], ensuring a fair comparison.
+Moreover, COLMAP fails on certain DTU scenes in the sparse-view setting.
 In these cases, we instead randomly initialize the point cloud.
@@
 […]
-3DGS [21] 18.52 0.523 0.415 0.167 22.80 0.708 0.276 0.096
-FSGS [53] 18.80 0.531 0.418 0.163 23.28 0.715 0.274 0.093
-CoR-GS [47] 19.52 0.558 0.418 0.151 23.39 0.727 0.271 0.091
-SE-GS 19.91 0.596 0.400 0.143 23.74 0.745 0.265 0.086
+3DGS [21] 18.52 0.523 0.415 22.80 0.708 0.276
+FSGS [54] 18.80 0.531 0.418 23.28 0.715 0.274
+CoR-GS [48] 19.52 0.558 0.418 23.39 0.727 0.271
+SE-GS 19.91 0.596 0.400 23.74 0.745 0.265
 Table 3: Results on Mip-NeRF360 with 12 and 24 training views. The first, second, and third-best results are marked in red, orange, and yellow, respectively.
@@
 In this context, our method surpasses all competitors across the four datasets.
-Note that in the sparse-view scenario, the available information in the training data is limited, making it more challenging to improve performance compared to the dense-view setting.
-Specifically, on the LLFF and DTU datasets, we train the methods using 3, 6, and 9 views.
-Our method achieves an improvement of $0.34$ in terms of PSNR with 3 views on LLFF.
+Note that in sparse-view scenarios, the available information in the training data is limited, making it more challenging to improve performance compared to the dense-view setting.
+Specifically, our method achieves an improvement of $0.34$ in terms of PSNR with 3 views on LLFF.
 The improvement is larger on DTU, reaching a gain of $0.59$.
 Since the images in the Mip-NeRF360 dataset are captured from diverse viewpoints, we sample more training views, i.e., 12 and 24, following the setting used in [48].
-On this dataset, our SE-GS achieves the highest PSNR and SSIM, as well as the lowest LPIPS and AVGE, demonstrating superior NVS quality under the challenges of sparse views with large-scale viewpoint variations.
+On this dataset, our SE-GS achieves the highest PSNR and SSIM, as well as the lowest LPIPS, demonstrating superior NVS quality under the challenges of sparse views with large-scale viewpoint variations.
 Additionally, to further assess the robustness across diverse scenarios, we conduct an experiment on 50 scenes sampled from the MVImgNet dataset.
@@
 However, the improvement is still not consistent, as CoR-GS shows worse LPIPS than 3DGS when trained with 10 views on the MVImgNet dataset.
-By contrast, our SE-GS beats 3DGS in all cases, demonstrating superior stability.
-Our method is capable of avoiding specific local optima as the $\mathbf{\Sigma}$-model aggregates information from diverse temporal samples in the presented self-ensembling mechanism.
+In contrast, our SE-GS beats 3DGS in all cases, demonstrating superior stability.
+This is attributed to its ability to effectively aggregate information from diverse perturbed models in the self-ensembling paradigm, leading to robust regularization.
 4.3 Qualitative Results
-Figure 5: Qualitative results. The methods are trained on sparse views and the renderings of novel views are illustrated. The images are selected from the DTU and MVImgNet datasets.
+Figure 5: Qualitative results. The methods are trained on sparse views and the renderings of novel views are illustrated. The images are from the DTU and MVImgNet datasets.
 Fig.
@@
 Our SE-GS produces fewer visual artifacts than the other methods, achieving better robustness against the challenge of few-shot training views.
-Moreover, our method captures finer details, particularly in areas with complex and repeated textures, such as the pineapple and watch strap.
+Moreover, our method captures finer details, particularly in areas with complex and repeated textures.
 This demonstrates the effectiveness of our self-ensembling strategy in improving both the visual quality and stability of novel view neural rendering under sparse-view conditions.
+Please refer to the supplementary material for more visualizations.
 (a) PSNR
@@
 The PSNR reaches a peak with 6 Gaussian models, indicating an upper bound of CoR-GS.
-Our method perturbs the $\mathbf{\Delta}$-model based on the uncertainties computed from dynamically updated image buffers, which results in more diverse samples of Gaussian models compared to CoR-GS.
+Our method perturbs the $\mathbf{\Delta}$-model based on the uncertainties computed from dynamically updated image buffers, which results in more diverse 3DGS models compared to CoR-GS.
 This leads to more effective regularization, as reflected by the better PSNR score.
@@
 This limitation poses a challenge in scaling CoR-GS up to a large number of Gaussian models.
-By contrast, in our pipeline, only two models are trained, i.e., the $\mathbf{\Delta}$-model and $\mathbf{\Sigma}$-model.
-The temporal samples are generated through perturbations.
-We achieve a comparable training speed to CoR-GS with 2 models, showing that the perturbation process incurs a negligible additional cost.
+In contrast, in our pipeline, only two models are trained, i.e., the $\mathbf{\Delta}$-model and $\mathbf{\Sigma}$-model.
+The perturbed models are generated through perturbation.
+In this context, we achieve a comparable training speed to CoR-GS with 2 models, showing that the perturbation process incurs a negligible additional cost.
 Note that only the $\mathbf{\Sigma}$-model is retained for NVS during inference, so the testing speed of SE-GS is the same as that of the original 3DGS.
@@
 Gradient-Aware denotes an alternative in which we perturb the Gaussians based on gradients.
-All methods are trained on LLFF with 3 views for 2000 iterations and the PSNR is utilized as the metric.
+All methods are trained on LLFF with 3 views for 2000 iterations, and PSNR is utilized as the metric.
+Figure 7: Impact of model diversity.
+The $x$-axis represents model agreement, indicating the similarity among models involved during training.
+Higher agreement values suggest greater model similarity and lower diversity.
+The $y$-axis denotes the PSNR of the methods during testing.
+The experiment is conducted on LLFF with three training views.
 Finally, we assess the effectiveness of our uncertainty-aware perturbation strategy, comparing it with two alternatives.
-Specifically, as introduced in Sec.
-3, a straightforward method for perturbing the $\mathbf{\Delta}$-model is to add random noise to the parameters of all Gaussians in the model.
+As introduced in Sec.
+3, a straightforward method for perturbing the $\mathbf{\Delta}$-model is to add random noise to the parameters of all Gaussians.
 We denote this approach as Random.
@@
 We train these two alternatives and our approaches on LLFF with 3 views for 2000 iterations, which challenges the methods in terms of convergence speed.
-As listed in Table 5, Random performs the worst since it leads to many invalid Gaussian samples, negatively impacting convergence speed.
-Note that our uncertainty-aware strategy aggregates information from multiple training steps, while the gradient-aware method can only utilize the gradients at the current training step.
-The improvement in PSNR shown in Table 5 demonstrates the superior effectiveness of our approach.
+As listed in Table 5, Random performs the worst, indicating that this strategy negatively impacts convergence speed.
+Notably, our uncertainty-aware strategy aggregates information from multiple training steps, while the gradient-aware method only utilizes the gradients at the current training step.
+The effectiveness of this design is evidenced by the improvement in PSNR shown in Table 5.
+Additionally, to further investigate the impact of model diversity, we conduct a detailed analysis on LLFF.
+As shown in Fig.
+7, the $x$-axis denotes the model agreement measured as PSNR between renderings of all involved models on the test views; the $y$-axis shows PSNR between the NVS result and the ground-truth image during testing.
+The Model agreement reflects consistency, where higher values indicate greater similarity among models.
+The agreements of CoR-GS with 5 and 8 models are nearly identical, meaning that the additional models lack sufficient diversity to enhance the ensembling process.
+This aligns with the observation in Fig.
+6(a) that merely using more models yields limited improvement.
+While perturbation can enhance model diversity, it may weaken supervision reliability.
+As shown by the red triangle in Fig.
+7, adding noise to all Gaussians yields overly diverse models.
+The supervision in the regularization term becomes too noisy, resulting in limited PSNR.
+In contrast, our uncertainty-aware perturbation method only perturbs unreliable Gaussians, enabling a good trade-off between model diversity and supervision reliability.
 5 Conclusion
 In this paper, we have presented a new self-ensembling mechanism that enhances the novel view synthesis of 3DGS with sparse training images.
-We have tackled the challenge of overfitting by training an ensemble named $\mathbf{\Sigma}$-model with the guidance of diverse temporal models sampled from the Gaussian parameter space.
-To obtain such samples, we have introduced an effective strategy that perturbs a $\mathbf{\Delta}$-model based on uncertainties computed from dynamically updated buffers of pseudo-view renderings.
+We have tackled the challenge of overfitting by training an ensemble named $\mathbf{\Sigma}$-model with the guidance of diverse perturbed models derived from a $\mathbf{\Delta}$-model.
+To obtain such models, we have introduced an effective strategy that perturbs the $\mathbf{\Delta}$-model based on uncertainties computed from dynamically updated buffers of pseudo-view renderings.
 We have conducted experiments on LLFF, DTU, Mip-NeRF360, and MVImgNet, as well as a comprehensive analysis of our approach.
 The experimental results have demonstrated the effectiveness, stability, and efficiency of our SE-GS.
-In future work, we plan to incorporate an outlier identifier into our perturbation mechanism to reduce the impact of unreliable samples in regularization and facilitate the self-ensembling process.
+In future work, we plan to incorporate an outlier identifier into our perturbation mechanism to reduce the impact of unreliable models in regularization and facilitate the self-ensembling process.
 Acknowledgment.
@@
 Supplementary Material
+Details on Ensembling Learning
+Figure 8: Pipeline of self-ensembling. Network variants are generated via dropout, and self-ensembling is achieved by utilizing a regularization term between the variants.
+Ensemble learning has been recognized for its effectiveness in improving robustness and generalization [13], especially in scenarios with limited training data.
+We illustrate the training pipeline of existing self-ensembling approaches [11] in Fig.
+8.
+Specifically, self-ensembling aims to aggregate information of diverse models without training multiple models from scratch.
+The model diversity is typically enhanced via dropout.
+In this context, the variants are dynamically derived from a single network during training, thereby prompting both diversity and efficiency.
+Self-ensembling is then achieved by incorporating a regularization term, which operates alongside the task-specific loss function based on ground truth.
+This regularization enforces consistency among the network variants by constraining their outputs, thereby mitigating model bias and enhancing robustness and generalization.
+Inspired by its effectiveness, we introduce the first self-ensembling Gaussian Splatting method to improve novel view synthesis in scenarios with sparse training views.
 Details on Uncertainty-Aware Perturbation
@@
 As shown in Fig.
-7, the curves represent the PSNR values obtained on testing views at different training iterations.
+9, the curves represent the PSNR values obtained on testing views at different training iterations.
 Perturbing the $\mathbf{\Delta}$-model too frequently, e.g., with an interval of 100, significantly degrades performance, indicating a negative impact on our method.
-Conversely, large intervals, such as 900, reduce the number of temporal samples, thereby weakening the self-ensembling effect and leading to decreased performance compared to the default setting.
+Conversely, large intervals, such as 900, reduce the number of perturbed models, thereby weakening the self-ensembling effect and leading to decreased performance compared to the default setting.
 Moreover, we analyze the effect of the noise level by conducting experiments with varying values of $\sigma$ in Eq.
@@
 The results on testing views at different training iterations are shown in Fig.
-8.
-Strong perturbations, such as those with $\omega$ values of 0.3 and 0.2, result in significantly worse performance compared to other settings.
-In these cases, the temporal models are sampled too far from the $\mathbf{\Delta}$-model in the Gaussian parameter space, reducing the consistency among temporal samples.
+10.
+Strong perturbation, such as those with $\omega$ values of 0.3 and 0.2, result in significantly worse performance compared to other settings.
+In these cases, the perturbed models are too far from the $\mathbf{\Delta}$-model in the Gaussian parameter space, reducing the consistency among the variants.
 The performance is also limited when a small $\omega$, such as 0.01, is used.
-In this scenario, all temporal models closely resemble the $\mathbf{\Delta}$-model, diminishing the benefits of self-ensembling.
+In this scenario, all perturbed models closely resemble the $\mathbf{\Delta}$-model, diminishing the benefits of self-ensembling.
 In contrast to using a fixed $\omega$, we adopt a decay function, where $\omega$ dynamically varies from 0.08 to 0.02 during training.
-This strategy achieves a good trade-off between the consistency and diversity of the temporal samples.
+This strategy achieves a good trade-off between the consistency and diversity of the perturbed models.
 Figure 11: Number of perturbed Gaussians throughout training. The curve illustrates the percentage of perturbed Gaussians in the $\mathbf{\Delta}$-model at different training iterations.
 To better understand the presented perturbation mechanism, in Fig.
-9, we illustrate the percentage of perturbed Gaussians in the $\mathbf{\Delta}$-model.
+11, we illustrate the percentage of perturbed Gaussians in the $\mathbf{\Delta}$-model.
 Note that the threshold $\tau$ in Eq.
@@
 Therefore, the percentage dynamically varies throughout training.
-This dynamic behavior balances the diversity and reliability of the temporal models, thereby enhancing the self-ensembling process.
+This dynamic behavior balances the diversity and reliability of the perturbed models, thereby enhancing the self-ensembling process.
 More Visualization Results
@@
 For more details, please refer to the uploaded videos.
-Notably, the videos of pseudo-view renderings and uncertainty maps showcase the results across different training iterations.
+nvs_res.mov: Shows the novel-view renderings of 3DGS, CoR-GS, and our SE-GS, where our method results in finer details and fewer artifacts.
+uncertainty_map_update.mov: Records the update of uncertainty maps during training.
+For each scene, we sample a pseudo view and store the corresponding rendering of the $\mathbf{\Delta}$-model in the image buffer.
+The uncertainty map is then derived from this buffer.
+The updates across different training iterations are shown in the video.
+sigma_model_vs_perturbed_model.mov: Displays the pseudo-view renderings of the $\mathbf{\Sigma}$-model and the perturbed $\mathbf{\Delta}$-model used in the regularization Eq.
+13.